### v1

# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

Abstract.
The dynamics of probability density functions has been extensively studied in science and engineering to understand physical phenomena and facilitate algorithmic design. Of particular interest are dynamics that can be formulated as gradient flows of energy functionals under the Wasserstein metric. The development of functional inequalities, such as the log-Sobolev inequality, plays a pivotal role in analyzing the convergence of these dynamics. The goal of this paper is to parallel the success of techniques using functional inequalities, for dynamics that are gradient flows under the Fisher-Rao metric, with various $f$-divergences as energy functionals. Such dynamics take the form of a nonlocal differential equation, for which existing analysis critically relies on using the explicit solution formula in special cases. We provide a comprehensive study on functional inequalities and the relevant geodesic convexity for Fisher-Rao gradient flows under minimal assumptions. A notable feature of the obtained functional inequalities is that they do not depend on the log-concavity or log-Sobolev constants of the target distribution. Consequently, the convergence rate of the dynamics (assuming well-posed) is uniform across general target distributions, making them potentially desirable dynamics for posterior sampling applications in Bayesian inference.
1. Introduction
The study of dynamics of probability density functions has proven invaluable across various disciplines, including statistical physics, developmental biology, and the design of efficient algorithms for uncertainty quantification. A particularly important application is sampling a target positive probability density ${\rho^{*}}:\mathbb{R}^{d}\to\mathbb{R}$ known up to normalization constants:
(1.1) ${\rho^{*}}(\theta)\propto\exp(-V(\theta)),$
in terms of a potential $V(\theta)$. For example, such sampling problems arise in climate science for calibrating climate models [46, 77, 45, 57], engineering for decision making and risk management [87, 26, 36, 14], tissue growth and cell populations models in biological engineering [42, 78, 34], and machine learning for quantifying prediction uncertainties [65, 41, 61, 86, 20]. The demand from these diverse applications has spurred extensive research into efficient and scalable sampling algorithms, a subject explored across applied mathematics, computer science, machine learning, and statistics.
Numerous approaches, based on constructing dynamics of probability density functions, have been proposed in the literature to address the sampling problem. The most widely used examples in Bayesian inference are sequential Monte Carlo (SMC) [29] and Markov chain Monte Carlo (MCMC) [13]. A classical instance in the MCMC category (in continuous time) is the Langevin dynamics [71]
(1.2) ${\rm d}X_{t}=\nabla\log\rho^{*}(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},$
and the corresponding Fokker-Planck equation
(1.3) $\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot(\rho_{t}\nabla\log\rho^{*})=\Delta\rho_{t},$
which satisfies that when $t\to\infty$, the law of $X_{t}$, namely $\rho_{t}$, converges to $\rho^{*}$, under certain regularity conditions on $\rho^{*}$. The study of the convergence rate has been of significant interest, as it characterizes the mixing time of the dynamics and the efficiency of the sampling algorithm [47, 23].
A remarkable observation is that the Fokker-Planck equation can be formulated as the gradient flow of the Kullback–Leibler (KL) divergence in information theory, known as the Boltzmann relative entropy in kinetic theory [81], with respect to the Wasserstein metric, according to the seminal work of Jordan, Kinderlehrer, and Otto [47]. For this reason, (1.3) is referred to as the Wasserstein gradient flow of the KL divergence. This gradient flow structure has profoundly influenced the analysis of the dynamics [15, 17, 18, 69] and the design of new sampling algorithms [84, 11, 21].
The goal of this paper is to systematically extend the success of the analysis to Fisher-Rao gradient flows which are nonlocal differential equations of great interest in statistics, information geometry and evolutionary game theory. These flows have the potential to achieve fast, universal convergence rates for sampling any target distribution, in contrast to Wasserstein gradient flows. In Section 1.1, we review the results for Wasserstein gradient flows, including functional inequalities (e.g., the log-Sobolev inequality), displacement convexity, and their induced convergence rates of the dynamics. Section 1.2 considers Fisher-Rao gradient flows and explains their significance in sampling applications and the difficulty in analyzing convergence. Section 1.3 presents a summary of our main results for a comprehensive study of geodesic convexity and functional inequalities in the Fisher-Rao setting. We demonstrate the favorable, uniform convergence rates of Fisher-Rao gradient flows across general target distributions. Related works are discussed in Section 1.4, and the organization of this paper is outlined in Section 1.5.
1.1. Wasserstein gradient flows: Functional inequalities and convexity
For the scope of this paper, we adopt the formal approach by Otto [68] to define gradient flows of probability densities. To do so, consider the probability space with smooth positive densities:
(1.4) $\mathcal{P}=\Bigl{\{}\rho\in C^{\infty}(\mathbb{R}^{d}):\int\rho(\theta)\mathrm{d}\theta=1,\,\rho>0\Bigr{\}}\,.$
The Riemannian metric at $\rho$ is denoted by $g_{\rho}$ such that
(1.5) $g_{\rho}(\sigma_{1},\sigma_{2})=\langle M(\rho)\sigma_{1},\sigma_{2}\rangle\ \text{for}\ \sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P},$
[…]
(1.6) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho_{t})=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr{|}_{\rho_{t}},$
where $\frac{\delta\mathcal{E}}{\delta\rho}$ is the first variation of the energy functional $\mathcal{E}$ such that
(1.7) $\Bigl{\langle}\frac{\delta\mathcal{E}}{\delta\rho},\sigma\Bigr{\rangle}=\lim_{\epsilon\rightarrow 0}\frac{\mathcal{E}(\rho+\epsilon\sigma)-\mathcal{E}(\rho)}{\epsilon},$
and we require $\mathbb{E}_{\rho}[\frac{\delta\mathcal{E}}{\delta\rho}]=0$ so that the first variation is unique.
By setting the energy function as the KL divergence
(1.8) $\mathcal{E}(\rho)=\text{KL}[\rho\|\rho^{*}]=\int\rho\log\frac{\rho}{\rho^{*}}{\rm d}\theta,$
and using the Wasserstein metric tensor $M(\rho)^{-1}\psi=-\nabla\cdot(\rho\nabla\psi)$, we obtain the Wasserstein gradient flow of the KL divergence:
(1.9) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho)=\nabla\cdot\Bigl{(}\rho_{t}\nabla\log\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)},$
which attains the same form as the Fokker-Planck equation (1.3).
The gradient flow structure of the Fokker-Planck equation naturally leads to Riemannian optimization interpretation for understanding its convergence [69]. In fact, along the gradient flow, it holds that
(1.10) $\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}(\rho_{t})-\mathcal{E}(\rho^{*}))=-g_{\rho_{t}}(\textrm{grad}_{M}\mathcal{E}(\rho_{t}),\textrm{grad}_{M}\mathcal{E}(\rho_{t})).$
If we have the following gradient dominance condition (also known as the Polyak-Łojasiewicz inequality [73]):
(1.11) $\exists\alpha>0,\ g_{\rho}(\textrm{grad}_{M}\mathcal{E}(\rho),\textrm{grad}_{M}\mathcal{E}(\rho))\geq 2\alpha(\mathcal{E}(\rho)-\mathcal{E}({\rho^{*}})),\textrm{ for any }\rho\in\mathcal{P},$
then, it is straightforward to derive $\mathcal{E}(\rho_{t})-\mathcal{E}({\rho^{*}})\leq\exp(-2\alpha t)\bigl{(}\mathcal{E}(\rho_{0})-\mathcal{E}({\rho^{*}})\bigr{)}$. In the case of $\mathcal{E}(\rho)=\text{KL}[\rho\|\rho^{*}]$ and $M(\rho)^{-1}\psi=-\nabla\cdot(\rho\nabla\psi)$, the gradient dominance condition is equivalent to
(1.12) $\exists\alpha>0\textrm{ s.t. }\int\rho\bigl{|}\nabla\log\frac{\rho}{{\rho^{*}}}\bigr{|}^{2}\mathrm{d}\theta\geq 2\alpha\int\rho\log\frac{\rho}{{\rho^{*}}}\mathrm{d}\theta,\textrm{ for any }\rho\in\mathcal{P}.$
This is precisely the log-Sobolev inequality [39, 38] for ${\rho^{*}}$ and $f=\sqrt{\frac{\rho}{{\rho^{*}}}}$:
(1.13) $\int\bigl{|}\nabla f(\theta)\bigr{|}^{2}{\rho^{*}}(\theta)\mathrm{d}\theta\geq\frac{\alpha}{2}\int f(\theta)^{2}\log\Bigl{(}\frac{f(\theta)^{2}}{\int f(\theta^{\prime})^{2}{\rho^{*}}(\theta^{\prime})\mathrm{d}\theta^{\prime}}\Bigr{)}{\rho^{*}}(\theta)\mathrm{d}\theta.$
An sufficient condition for the gradient dominance condition is the strong geodesic convexity:
(1.14) $\exists\alpha>0,\ g_{\rho}(\textrm{Hess}_{M}\mathcal{E}(\rho)\sigma,\sigma)\geq\alpha g_{\rho}(\sigma,\sigma),\textrm{ for any }\rho\in\mathcal{P},\sigma\in T_{\rho}\mathcal{P},$
where $\textrm{Hess}_{M}\mathcal{E}(\rho)$ is the Riemannian Hessian of $\mathcal{E}$. For the Wasserstein case $M(\rho)^{-1}\psi=-\nabla\cdot(\rho\nabla\psi)$, the geodesic convexity is also known as the displacement convexity [64]. The Bakry-Emery theorem [8, 9] establishes the connection between the displacement convexity, the log-concavity of $\log\rho^{\star}$, and the Ricci curvature of the sample space, and conditions for general convex Sobolev inequalities involving different relative entropies or divergences were established in [81, 5]. In summary, these functional inequalities and convexity properties play important roles for studying the convergence of dynamics arising from Wasserstein gradient flows [64, 17, 3].
We note that while the above explanations regarding the geometry and gradient flow structures are mostly formal, they can be made rigorous [3, 76]. On the other hand, it is also quite important to use strategically the geometric structure to formally derive the sought inequalities. Then, prove rigorously the resulting inequalities to eventually show the convergence of the dynamics once well-posedness is analysed. In this work, we focus on rigorously proving the functional inequalities.
1.2. Analysis of Fisher-Rao gradient flows
In this paper, our main focus is on the gradient flow dynamics with respect to the Fisher-Rao metric [2, 7], which takes the following form
(1.15) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
The corresponding metric tensor is $M^{\mathrm{FR}}(\rho)^{-1}\psi=\psi-\mathbb{E}_{\rho}[\psi]$. As an example, the Fisher-Rao gradient flow of the KL-divergence can be expressed as follows:
(1.16) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{\rm FR}{\rm KL}[\rho_{t}\|{\rho^{*}}]=-\rho_{t}\Bigl{(}\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)}.$
Recently, convergence of the above Fisher-Rao gradient flow (1.16), namely
(1.17) ${\rm KL}[\rho_{t}\|{\rho^{*}}]=\mathcal{O}(\exp(-t)).$
has been established in [59, 21, 60, 28]. We highlight that the exponential convergence rate here is uniform and independent of $\rho^{*}$. This characteristic stands in sharp contrast to the Wasserstein gradient flow (1.9), where the convergence rate $\mathcal{O}(\exp(-\alpha t))$ depends on the log-Sobolev constant $\alpha$; this constant depends on $\rho^{*}$ and can be very small for anisotropic or multimodal distributions. As a consequence, the uniform convergence rate of the Fisher-Rao gradient flow makes it a potentially desirable dynamics for sampling applications; see [59, 60, 44, 56, 22, 63].
Nevertheless, all the current proofs of the uniform convergence of the Fisher-Rao gradient flow (1.16) rely on the explicit formula of the solution
(1.18) $\rho_{t}(\theta)\propto\rho_{0}(\theta)^{e^{-t}}{\rho^{*}}(\theta)^{1-e^{-t}},$
which is specific to the KL divergence as the energy functional. Moreover, the convergence results require strong assumptions on $\rho_{0}$. This prevents us from extending the convergence framework of Fisher-Rao gradient flows to general energy functionals and initial conditions $\rho_{0}$. On the other hand, the Fisher-Rao gradient flow, although having the form of an ODE at each $\theta$, contains a nonlocal interaction term that conserves the mass: $\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log\rho^{*}]$. It is non-trivial to analyze such nonlocal terms without using the explicit formula.
In this paper, we address the above challenge by providing a systematic study of functional inequalities and geodesic convexity for the Fisher-Rao gradient flows with general $f$-divergences as energy functionals. Here, the $f$-divergences are defined as
(1.19) $D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta,$
where $f(1)=0,f\in C^{2}(0,+\infty)$ and $f$ is convex. It is worthy noting that $f$-divergences are known as relative entropies in the context of statistical physics and kinetic theory, see for instance [81, 5, 4].
1.3. Main results
Firstly, we show that the gradient dominance condition and the geodesic convexity do not hold for the Fisher-Rao gradient flows of the KL divergence (1.16). This poses significant challenges for establishing the convergence of the dynamics using the functional inequality approach. The informal statement is presented below and the detailed analysis is in Theorems 3.3 and 4.1.
Theorem 1.1 (informal).
Under the Fisher-Rao geometry, the KL divergence is not geodesically convex for any ${\rho^{*}}$, even within a small neighborhood of ${\rho^{*}}$. Moreover, the gradient dominance condition is not satisfied for any ${\rho^{*}}$, also within any small neighborhood of ${\rho^{*}}$. Here the “small neighborhood” is measured in the KL divergence.
Secondly, for general $f$-divergences as the energy functionals, we derive a sufficient condition for geodesic convexity under the Fisher-Rao geometry, and a sufficient and necessary condition for the gradient dominance condition to hold. These lead to general convergence results for the corresponding Fisher-Rao gradient flows. We state the informal result below and the detailed result is in Theorems 3.4, 4.2 and 4.6; our proofs use novel techniques from linear programming in optimization to reduce the challenging global nonlocal interaction to a three-point interaction which is much simpler to deal with.
Theorem 1.2 (informal).
Under the Fisher-Rao geometry, the $f$-divergence is geodesically convex when $xf^{\prime}(x)$ is concave. Additionally under the condition $f^{\prime\prime}(1)>0$, the $f$-divergence achieves $\alpha_{f}$-strong geodesic convexity with $\alpha_{f}>0$ depending solely on $f$.
Moreover, for general $f$-divergences, the gradient dominance condition holds with an $\alpha_{f}>0$ depending solely on $f$ if and only if
$\begin{split}\exists\alpha>0\qquad\big{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha\left(\frac{1}{x}-\frac{1}{y}\right)\left(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\right)\quad\text{for all }\ 0<x<1<y.\end{split}$
An sufficient condition is that there exists a constant $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Note that the KL divergence does not satisfy the conditions.
Thirdly, we study the case of the KL divergence in further depth and seek other possible functional inequalities to establish the convergence. We prove a novel dual type of gradient dominance conditions that allow us to obtain the convergence of the Fisher-Rao gradient flows for general $f$-divergences. To the best of our knowledge, such dual type of functional inequalities is not known in the Wasserstein case and it is a nontrivial, unique discovery in the Fisher-Rao case. The informal statement is as follows; the detailed statement is in Section 5.
Theorem 1.3 (informal).
For the Fisher-Rao gradient flow $\rho_{t}$ of an $f$-divergence, we consider its evolution under another $\bar{f}$-divergence. We define the dual gradient dominance condition on this flow as
(1.20) $\frac{\rm d}{{\rm d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]\leq-\alpha_{f}\Bigl{(}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]+D_{f}[\rho_{t}\|{\rho^{*}}]\Bigr{)}.$
For $f\in C^{2}(0,+\infty)$ and $f^{\prime\prime}(1)>0$, the above inequality holds with an $\alpha_{f}$ independent of $\rho^{*}$, if we set ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$.
Furthermore, for the KL divergence, if we take ${\bar{f}}(x)=xf(\frac{1}{x})$, then the dual gradient dominance condition holds with an explicit $\alpha_{f}=1$. This can also be generalized to $f$-divergences with $f$ being a polynomial.
With these established functional inequalities, it is straightforward to show the corresponding gradient flows converge to the target exponentially fast at the same rate, given that the gradient flows are well posed, allowing us to differentiate the energy functional along the dynamics. The main focus of this paper is the rigorous proof of the functional inequalities and we leave the study of the well-posedness of the gradient flows for future work.
1.4. Related works
1.4.1. Displacement convexity and functional inequalities
With respect to the Wasserstein metric, geodesic convexity is equivalent to the displacement convexity, a concept introduced by McCann in [64]. Further exploration of displacement convexity on Riemannian manifolds and more general spaces can be found in [69, 25, 70, 82, 58, 66]. Additionally, the work [16] studies the displacement convexity concerning the generalized Wasserstein metric with nonlinear mobility. These concepts have been extended to more general settings such as graphs, discrete space, Markov chains with other metrics, see for instance [31, 32, 30].
Various functional inequalities related to the Wasserstein metric have been developed, such as the Poincaré inequality, log-Sobolev inequality, and Talagrand’s inequality; see [72, 80, 69]. These functional inequalities can be used to study the convergence of Wasserstein gradient flows and related PDEs [69]. Log-Sobolev type inequalities have been proved for general divergences or relative entropy and for general nonlinear and nonlocal McKean-Vlasov equations; see for instance [5, 15, 17, 3]. However, few functional inequalities have been developed for the Fisher-Rao gradient flow.
1.4.2. Fisher-Rao gradient flow
The Fisher-Rao metric and its finite-dimensional counterpart in parameter space, the Fisher information matrix, are fundamental concepts in information geometry [2, 7]; see also [74, 35, 79]. The geodesic distance under the Fisher-Rao metric is known as the Fisher-Rao distance, which, under a mass conservation constraint, is also referred to as the spherical Hellinger distance [53, 52, 12]. The dynamical formulation of the Fisher-Rao distance is related to the development of unbalanced optimal transport and the Wasserstein-Fisher-Rao distance [24, 51, 54].
The Fisher-Rao gradient flow of the KL divergence (1.16) is a mean-field equation of birth-death type [59, 60]:
$\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Lambda_{t},\qquad\Lambda_{t}=\bigl{(}\log\rho_{t}-\log{\rho^{*}}\bigr{)}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}],$
where the birth-death rate $\Lambda_{t}$ contains a nonlocal interaction term that allows global movement of mass. Similar equations in discrete space, known as the replicator equations, have been studied in evolutionary game theory [43]. Projection of the Fisher-Rao gradient flows to a parametric family of distributions leads to natural gradient flows. Time discretizations of these flows, known as natural gradient descent algorithms, have been widely used in variational inference [1, 55, 62, 21].
Fisher-Rao gradient flows have found applications in various fields. For example, the corresponding birth-death process has been used in sequential Monte Carlo samplers to reduce the variance of particle weights [27], and in the acceleration of Langevin sampling [59, 60] and training of neural networks [75]. The Fisher-Rao gradient flow of the $\chi^{2}$ distance arises as a mean-field equation for ensemble MCMC methods that address multimodal distributions [56]. Kernel approximations of Fisher-Rao gradient flows have been proposed for sampling [63]. Furthermore, parametric approximation of Fisher-Rao gradient flows employing Kalman’s methodology [49, 33, 48] has led to the development of efficient, derivative-free posterior approximation algorithms. These algorithms, based on Gaussian or Gaussian mixture approximations [44, 22], have proven particularly useful for Bayesian inverse problems. Other applications in filtering [40] and reinforcement learning [50] have also been explored in the literature.
1.5. Organization of this paper
In Section 2, we introduce Fisher-Rao gradient flows associated with various $f$-divergence. In Section 3, we present an overview of convex analysis and study the geodesic convexity of the $f$-divergence in the Fisher-Rao geometry. In Section 4, we examine the gradient dominance condition of $f$-divergence under the Fisher-Rao metric. In Section 5, we develop the dual gradient dominance condition for analyzing the convergence properties of $f$-divergences that do not satisfy the gradient dominance condition. We conclude in Section 6.
2. Fisher-Rao Gradient Flows of $f$-divergences
We follow the approach in Section 1.1 to derive the equation of Fisher-Rao gradient flows of $f$-divergences. Again, the derivation is formal, and our main goal is to analyze the relevant functional inequalities arising from such gradient flows, which allow us to establish convergence rates of the dynamics when they are well-posed.
We consider the energy functional to be the $f$-divergence
(2.1) $\mathcal{E}(\rho)=D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta,$
where $f(1)=0,f\in C^{2}(0,+\infty)$ and $f$ is convex. Here, the convexity of $f$ implies
$f(\rho/{\rho^{*}})\geq f(1)+f^{\prime}(1)(\rho/{\rho^{*}}-1),$
which leads to
(2.2) $D_{f}[\rho\|{\rho^{*}}]\geq\int{\rho^{*}}\Bigl{(}f(1)+f^{\prime}(1)(\rho/{\rho^{*}}-1)\Bigr{)}\,\mathrm{d}\theta\geq-\int\bigl{|}f^{\prime}(1)(\rho-{\rho^{*}})\bigr{|}\,\mathrm{d}\theta\geq-2\lvert f^{\prime}(1)\rvert.$
This means $f$-divergence is always well defined (with a possible value +$\infty$+ ∞). Moreover, Jensen’s inequality implies that ${\rho^{*}}$ is the unique minimizer. The KL-divergence is a special $f$-divergence with $f=x\log x$.
For the $f$-divergence, its first variation with respect to $\rho$ is $f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}[f^{\prime}(\frac{\rho}{{\rho^{*}}})]$ where we subtract a constant so that its mean is zero. Using the formula of the Fisher-Rao metric (1.15), we get the corresponding Fisher-Rao gradient flow as
(2.3) $\frac{\partial\rho_{t}}{\partial t}=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr{|}_{\rho_{t}}=-\rho_{t}f^{\prime}\Bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)}+\rho_{t}\mathbb{E}_{\rho_{t}}\Bigl{[}f^{\prime}\Bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)}\Bigr{]}.$
We note that the Fisher-Rao metric (1.15) stands out as the unique metric, up to constants, that satisfies the diffeomorphism invariance property [19, 6, 10]. Under the Fisher-Rao metric, these gradient flows are expected to maintain this invariance. The diffeomorphism invariance implies that the convergence property of Fisher-Rao gradient flows for a general target density ${\rho^{*}}$ aligns with that for the standard Gaussian target density, and may achieve a convergence rate independent of ${\rho^{*}}$. For this reason, they have been widely used in sampling applications.
When $f(x)=x\log x$ which corresponds to the KL divergence, we get
(2.4) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)},$
which is the most commonly used Fisher-Rao gradient flow for sampling applications [59, 21, 60, 28, 85]. When $f(x)=(x-1)^{2}$ which corresponds to the $\chi^{2}$ divergence, we get
(2.5) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\frac{\rho_{t}}{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\frac{\rho_{t}}{\rho^{*}}]\Bigr{)},$
which has been used in the context of ensemble samplers [56].
3. Geodesic Convexity in the Fisher-Rao Geometry
In this section, we explore the geodesic convexity of $f$-divergences in the Fisher-Rao geometry. We will present counterexamples to demonstrate that the KL divergence lacks geodesic convexity. Subsequently, we will provide sufficient conditions to identify the specific $f$-divergences that achieve geodesic convexity. The discussions in this section parallel those of displacement convexity in the Wasserstein geometry. The geometric calculations are primarily carried out in a formal basis  [^1] However, we note that they can be made rigorous by restricting to probability simplex $\sum_{i=1}^{K}\rho_{i}=1,\rho_{i}>0$, where the Riemannian geometry is rigorously defined.. We first outline the preliminary concepts in the Euclidean space and abstract manifolds and then move to the Fisher-Rao geometry.
3.1. Preliminaries on convex analysis
We start with introducing the basic concepts in the Euclidean space. Let $h$ be a convex function in $\mathbb{R}^{N}$, then for any $\theta_{0},\theta_{1}\in\mathbb{R}^{N}$ and $\lambda\in[0,1]$, it holds that
(3.1) $h(\lambda\theta_{0}+(1-\lambda)\theta_{1})\leq\lambda h(\theta_{0})+(1-\lambda)h(\theta_{1}).$
Furthermore suppose $h\in C^{2}(\mathbb{R}^{N})$, then we have the first order characterization that
(3.2) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle,$
and the second order characterization that the Hessian matrix $\nabla^{2}_{\theta}h(\theta)\succeq\alpha I$ for some $\alpha\geq 0$ and any $\theta\in\mathbb{R}^{N}$. When $\alpha>0$, we call the function $h$ to be $\alpha$-strongly convex; in such case we have
(3.3) $h(\theta_{1})\geq h(\theta_{0})+\langle\nabla_{\theta}h(\theta_{0}),\theta_{1}-\theta_{0}\rangle+\frac{\alpha}{2}|\theta_{1}-\theta_{0}|^{2}.$
For a gradient flow in the Euclidean space, denoted by
(3.4) $\frac{{\rm d}\theta_{t}}{{\rm d}t}=-\nabla_{\theta}h(\theta_{t}),$
the evolution of the function values satisfy
(3.5) $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})=-|\nabla_{\theta}h(\theta_{t})|^{2}.$
Equation 3.3 indicates that $h$ has a unique global minimizer, denoted by $\theta^{*}$. Using the strong convexity of $h$ we have
(3.6) $h(\theta)-h({\theta^{*}})\leq\langle\nabla_{\theta}h(\theta),\theta-{\theta^{*}}\rangle-\frac{\alpha}{2}|{\theta^{*}}-\theta|^{2}\leq\frac{1}{2\alpha}|\nabla_{\theta}h(\theta)|^{2},$
where in the last inequality we used the fact that $\langle a,b\rangle\leq\frac{1}{2\alpha}|a|^{2}+\frac{\alpha}{2}|b|^{2}$. Therefore, we obtain
(3.7) $|\nabla_{\theta}h(\theta)|^{2}\geq 2\alpha(h(\theta)-h({\theta^{*}})),$
which is referred to as the gradient dominance condition or the Polyak-Łojasiewicz inequality [73]. With this inequality, the gradient flow converges exponentially fast, as $\frac{{\rm d}}{{\rm d}t}h(\theta_{t})\leq-2\alpha(h(\theta_{t})-h({\theta^{*}}))$ leading to $h(\theta_{t})-h({\theta^{*}})=\mathcal{O}(\exp(-2\alpha t))$. In a summary, the strong convexity in the Euclidean space leads to gradient dominance condition, which then implies the exponential convergence of the gradient flow.
Now, consider a Riemannian manifold $\mathcal{M}$. For $p\in\mathcal{M}$, the Riemannian metric is $g_{p}:T_{p}\mathcal{M}\times T_{p}\mathcal{M}\to\mathbb{R}$. Geodesics in the manifold generalizes the concept of straight lines in Euclidean spaces. A constant-speed geodesics connecting two points $p_{0},p_{1}\in\mathcal{M}$ can be characterized as the minimizer of the following variational problem:
(3.8) $\mathcal{D}^{2}(p_{0},p_{1})=\inf_{p_{t}\in\mathcal{M},v_{t}\in T_{p}\mathcal{M}}\{\int_{0}^{1}g_{p_{t}}(v_{t},v_{t}){\rm d}t:\ \frac{{\rm d}}{{\rm d}t}p_{t}=v_{t},p_{t=0}=p_{0},p_{t=1}=p_{1}\},$
where $\mathcal{D}$ is called the geodesic distance between $p_{0}$ and $p_{1}$. Such a geodesics satisfies $g_{p_{t}}(v_{t},v_{t})=\mathcal{D}^{2}(p_{0},p_{1})$ so is referred to as the constant-speed geodesics.
A function $h:\mathcal{M}\to\mathbb{R}$ is called geodesically convex, if for any $p_{0},p_{1}\in\mathcal{M}$, $\lambda\in[0,1]$, it holds that
(3.9) $h(p_{\lambda})\leq\lambda h(p_{0})+(1-\lambda)h(p_{1}),$
where $p_{t},t\in[0,1]$ is the constant-speed geodesics connecting $p_{0}$ and $p_{1}$. The first order characterization of geodesic convexity is
(3.10) $h(p_{1})\geq h(p_{0})+g_{p_{0}}(\text{grad}~{}h(p_{0}),v_{0})+\frac{\alpha}{2}\mathcal{D}^{2}(p_{0},p_{1}),$
where grad is the gradient operator on the manifold, $v_{0}$ is the initial velocity of the constant-speed geodesics connecting $p_{0}$ and $p_{1}$. Here $h$ is called $\alpha$-strongly geodesically convex if $\alpha>0$. Furthermore, the second order characterization of the geodesic convexity is $\text{Hess}~{}h(p_{0})\succeq\alpha I$ where Hess is the Hessian operator defined on the manifold, which satisfies
(3.11) $g_{p_{0}}(\textrm{Hess}\,h(p_{0})v_{0},v_{0})=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big{|}_{\rm geod}h(p_{0})$
where $p_{t}$ is a constant-speed geodesics that starts at $p_{0}$ with initial velocity vector $v_{0}$. Here we used the notation $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big{|}_{\rm geod}$ to specifically mean that the derivative is taken with respect to a constant geodesics $p_{t}$ at $t=0$.
Similarly as the Euclidean case, $\alpha$-strongly geodesic convexity implies the gradient dominance condition:
(3.12) $g_{p}(\text{grad}\,h(p),\text{grad}\,h(p))\geq 2\alpha(h(p)-h(p^{*})),$
where $p^{*}$ is the global minimizer of $h$. Again, using this inequality leads to the exponential convergence of gradient flows of $h$ on the manifold $\mathcal{M}$.
3.2. Hessian operator in the Fisher-Rao geometry
In this subsection, we derive the Hessian operator of energy functionals in the Fisher-Rao geometry, which will be used in the next subsection to analyze the geodesic convexity of $f$-divergences.
The Fisher-Rao metric in the probability density space $\mathcal{P}$ is
(3.13) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
The geodesic distance between $\rho_{1},\rho_{1}$, referred to as the Fisher-Rao distance, satisfies
(3.14a) $\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})$ $=\inf_{\rho_{t},\sigma_{t}}\Bigl{\{}\int_{0}^{1}\int\frac{|\sigma_{t}|^{2}}{\rho_{t}}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\sigma_{t}\Bigr{\}}$
(3.14b) $=\inf_{\rho_{t},\psi_{t}}\Bigl{\{}\int_{0}^{1}\int\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])^{2}\mathrm{d}\theta\mathrm{d}t:\partial_{t}\rho_{t}=\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])\Bigr{\}}.$
Remark 3.1.
The Fisher-Rao geodesic distance attains an explicit formula, equivalent to (up to a constant scaling) the spherical Hellinger distance:
$\mathcal{D}_{\rm FR}^{2}(\rho_{0},\rho_{1})=4\operatorname{arccos}^{2}\left(\int\sqrt{\rho_{0}}\sqrt{\rho_{1}}{\rm d}\theta\right).$
The constant-speed geodesics is
$\rho_{t}(\theta)=\frac{\left(t\sqrt{\rho_{0}(\theta)}+(1-t)\sqrt{\rho_{1}(\theta)}\right)^{2}}{(1-t)^{2}+t^{2}+2t(1-t)\int\sqrt{\rho_{0}(\theta^{\prime})\rho_{1}(\theta^{\prime})}{\rm d}\theta^{\prime}}.$
For more discussions, see [40, 53, 60]. In view of this relation, Fisher-Rao gradient flows are sometimes referred to as spherical Hellinger gradient flows in the literature [56, 60].
$\lozenge$
Remark 3.2.
If we do not restrict the distributions $\rho_{t}$ to be on the probability space and we allow them to have any positive mass, then by using the relation
$\frac{|\sigma_{t}|^{2}}{\rho_{t}}=\frac{|\partial_{t}{\rho}_{t}|^{2}}{\rho_{t}}=4\Bigl{|}\frac{\mathrm{d}}{\mathrm{d}t}\sqrt{\rho_{t}}\Bigr{|}^{2}$
and the Cauchy-Schwarz inequality, we will get that the optimal objective value in (3.14) is $4\int|\sqrt{\rho_{0}}-\sqrt{\rho_{1}}|^{2}\mathrm{d}\theta.$ This is (up to a constant scaling) the Hellinger distance [37]. $\lozenge$
Using a formal Lagrangian duality calculation, we can derive the optimality condition of the variational problem (3.14) as a coupled system of differential equations (see also [83]):
(3.15a) $\partial_{t}\rho_{t}$ $=\rho_{t}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}]),$
(3.15b) $\partial_{t}\psi_{t}$ $=-\frac{1}{2}(\psi_{t}-\mathbb{E}_{\rho_{t}}[\psi_{t}])^{2},$
along with boundary condition $\rho_{t=0}=\rho_{0},\rho_{t=1}=\rho_{1}$. This can be utilized to calculate the gradient and Hessian operators.
Consider any constant-speed geodesics starting at $\rho$ with initial velocity field $\sigma=\rho(\psi-\mathbb{E}_{\rho}[\psi])$. Differentiating along this geodesics leads to the formula of the gradient and Hessian operator in the Fisher-Rao geometry:
(3.16a) $g_{\rho}^{\rm FR}(\textrm{grad}_{\rm FR}\mathcal{E}(\rho),\sigma)=$ $\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{\rm geod}\mathcal{E}(\rho)=\int\frac{\delta\mathcal{E}}{\delta\rho}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta,$
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}\mathcal{E}(\rho)\sigma,\sigma)=$ $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\Big{|}_{\rm geod}\mathcal{E}(\rho)=\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{\rm geod}\int\frac{\delta\mathcal{E}}{\delta\rho}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta$
(3.16b) $=$ $\int\frac{\delta^{2}\mathcal{E}}{\delta\rho^{2}}\rho^{2}(\psi-\mathbb{E}_{\rho}[\psi])^{2}+\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}\frac{\delta\mathcal{E}}{\delta\rho}-\mathbb{E}_{\rho}[\frac{\delta\mathcal{E}}{\delta\rho}]\Bigr{)}\mathrm{d}\theta.$
3.3. Geodesic convexity in the Fisher-Rao geometry
For the $f$-divergences (2.1), the corresponding gradient and Hessian operators are
(3.17a) $g_{\rho}^{\rm FR}(\textrm{grad}_{\rm FR}D_{f}[\rho\|{\rho^{*}}],\sigma)=$ $\int f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\rho(\psi-\mathbb{E}_{\rho}[\psi])\mathrm{d}\theta,$
(3.17b) $g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)=$ $\int f^{\prime\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\frac{\rho^{2}(\psi-\mathbb{E}_{\rho}[\psi])^{2}}{{\rho^{*}}}$
$+$ $\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}[f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}]\Bigr{)}\mathrm{d}\theta,$
where $\sigma=\rho(\psi-\mathbb{E}_{\rho}[\psi])$.
We employ the aforementioned formula to investigate the geodesic convexity of $f$-divergences within the Fisher-Rao geometry. In what follows, we first construct counterexamples to demonstrate that the KL divergence fails to exhibit geodesic convexity (see Theorem 3.3). Subsequently, we provide a sufficient condition for an $f$-divergence to satisfy geodesic convexity (see Theorem 3.4).
Theorem 3.3.
For any smooth, positive target density $\rho^{*}$, the KL divergence $\mathrm{KL}[\rho\|\rho^{*}]$ is not geodesically convex within the Fisher-Rao geometry.
Furthermore, for any $\epsilon>0$, the KL divergence is not geodesically convex in the neighborhood of $\rho^{*}$ defined by $\mathrm{KL}[\rho\|\rho^{*}]\leq\epsilon$.
Proof.
The KL-divergence corresponds to the $f$-divergence with $f(x)=x\log x$. In such case, the Hessian formula (3.17b) has the form
(3.18) $\begin{split}g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}{\rm KL}[\rho\|{\rho^{*}}]\sigma,\sigma)&=\int\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\\ &+\frac{1}{2}\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}\mathrm{d}\theta.\end{split}$
Thus, whether the Hessian is positive semi-definite or not depends on the sign of $1+\frac{1}{2}(\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}])$. The KL divergence will be geodesically convex when this term is non-negative. A simple sufficient condition is $1/e\leq{\rho}/{{\rho^{*}}}\leq e$.
However, in general, this term does not have a definite sign, and the KL divergence is not geodesically convex. To show that the Hessian is not positive semi-definite, we set a specific $\psi=\log\frac{\rho}{{\rho^{*}}}$, then (3.18) becomes
(3.19) $\mathcal{H}(\rho,{\rho^{*}})=\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{2}+\frac{1}{2}\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{3}\mathrm{d}\theta.$
We construct the following two examples to explain why the above term is not non-negative in general.
We first consider 1D Gaussian densities $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma^{2})$ and ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$. We have that
(3.20) $\begin{split}&\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]=\frac{\sigma^{2}}{2}+\frac{\mu^{2}}{2}-\log\sigma-\frac{1}{2},\\ &\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{2}\mathrm{d}\theta=\frac{1}{2}(\sigma^{2}-1)^{2}+\sigma^{2}\mu^{2},\\ &\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-\mathbb{E}_{\rho}[\log\frac{\rho}{{\rho^{*}}}]\Bigr{)}^{3}\mathrm{d}\theta=(\sigma^{2}-1)^{3}+3(\sigma^{2}-1)\sigma^{2}\mu^{2}.\end{split}$
Plugging (3.20) into (3.19) leads to that
$\mathcal{H}(\rho,{\rho^{*}})=\frac{1}{2}\sigma^{2}(\sigma^{2}-1)^{2}+\frac{3}{2}\sigma^{4}\mu^{2}-\frac{1}{2}\sigma^{2}\mu^{2}=\frac{\sigma^{2}}{2}\bigl{(}(\sigma^{2}-1)^{2}+3\sigma^{2}\mu^{2}-\mu^{2}\bigr{)}.$
Therefore, when
$\sigma^{2}<\frac{1}{3}\qquad\textrm{and}\qquad\mu^{2}>\frac{(\sigma^{2}-1)^{2}}{1-3\sigma^{2}},$
we have that $\mathcal{H}(\rho,{\rho^{*}})<0$.
Next, we discuss a general example beyond the Gaussian case. We will first give a counterexample where the density is not continuous. We then discuss the construction of a counterexample with a smooth density, by convolving the previous example with a smooth bump function.
Given any ${\rho^{*}}$ and radius $r>0$, we denote
(3.21) $\rho_{r}:=\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta,$
and construct the following density
(3.22) $\rho(\theta)={\rho^{*}}(\theta)(x_{1}\mathbbm{1}_{S^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{S_{r}}(\theta))$
where $\mathbbm{1}_{A}$ is the indicator function of a set $A$; $S_{r}$ denotes the sphere with radius $r$. For $\rho(\theta)$ to be a probability density, it is necessary that
(3.23) $x_{1}\Bigl{(}1-\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta\Bigr{)}+x_{2}\int_{S_{r}}{\rho^{*}}(\theta)\mathrm{d}\theta=x_{1}(1-\rho_{r})+x_{2}\rho_{r}=1.$
To see why this is the case, note that we can vary $r\in(0,+\infty)$ such that $\rho_{r}$ in (3.21) can take any value in $(0,1)$. Thus, for any choice of $x_{1}>1>x_{2}>0$, we can take some $r\in(0,+\infty)$ such that
$\rho_{r}=\frac{x_{1}-1}{x_{1}-x_{2}}\in(0,1),$
which leads to (3.23).
We can compute KL$[\rho\|{\rho^{*}}]$ and $\mathcal{H}(\rho,{\rho^{*}})$ denoted in (3.19), as follows:
${\rm KL}[\rho\|{\rho^{*}}]=$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}+\frac{x_{1}x_{2}-x_{2}}{x_{1}-x_{2}}\log x_{2},$
$\mathcal{H}(\rho,{\rho^{*}})=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}+\frac{1}{2}\rho_{r}(1-\rho_{r})x_{1}x_{2}\bigl{(}x_{2}\rho_{r}-x_{1}(1-\rho_{r})\bigr{)}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{3}$
$=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1+\frac{x_{2}\rho_{r}-x_{1}(1-\rho_{r})}{2}\log\frac{x_{1}}{x_{2}}\Bigr{)}$
$=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1-\frac{x_{1}+x_{2}-2x_{1}x_{2}}{2(x_{1}-x_{2})}\log\frac{x_{1}}{x_{2}}\Bigr{)}.$
Let us choose $x_{1}=e^{\epsilon}$ with $\epsilon>0$ and $x_{2}=e^{-M}$ with $M>4$. Then we have
(3.24a) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon,$
(3.24b) $\mathcal{H}(\rho,{\rho^{*}})=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1-\frac{x_{1}+x_{2}-2x_{1}x_{2}}{2(x_{1}-x_{2})}(M+\epsilon)\Bigr{)}$
$=$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1-\frac{e^{M}+e^{-\epsilon}-2}{2(e^{M}-e^{-\epsilon})}(M+\epsilon)\Bigr{)}$
$\leq$ $\rho_{r}(1-\rho_{r})x_{1}x_{2}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}\Bigl{(}1-\frac{e^{M}-2}{2e^{M}}M\Bigr{)}<0.$
Finally, to smooth $\rho$ in (3.22), we introduce the mollifier:
$\phi_{\delta}(\theta)=\begin{cases}\frac{1}{Z_{\delta}}e^{-\frac{1}{1-|\theta/\delta|^{2}}}&|\theta|<\delta\\ 0&|\theta|\geq\delta,\end{cases}$
which is smooth, has a compact support within $B_{\delta}$, and satisfies $\int\phi_{\delta}(\theta)\mathrm{d}\theta=1$. The smoothed density is constructed using convolution as follows:
$\rho_{\delta}(\theta)={\rho^{*}}(\theta)\bigl{(}\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{3}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}})\bigr{)}(\theta).$
Here, $x_{3}$ is determined by $\int\rho_{\delta}(\theta)\mathrm{d}\theta=1$, and it satisfies $x_{2}\leq x_{3}\leq x_{1}$, since
$\begin{split}\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{2}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}})\leq&x_{1}\mathbbm{1}_{S^{c}_{r}}+x_{2}\mathbbm{1}_{S_{r}}\\ \leq&\phi_{\delta}*(x_{1}\mathbbm{1}_{S^{c}_{r+\delta}}+x_{2}\mathbbm{1}_{S_{r-\delta}}+x_{1}\mathbbm{1}_{S_{r+\delta}\backslash S_{r-\delta}}).\end{split}$
This also leads to that $x_{2}\leq\frac{\rho_{\delta}}{{\rho^{*}}}\leq x_{1}$. Since $\rho_{\delta}(\theta)$ only deviates from $\rho$ (as denoted in (3.22)) within the compact band $S_{r+2\delta}\backslash S_{r-2\delta}$, with the measure depending on $\delta$, the estimations in eq. 3.24 remain valid when $\delta$ is sufficiently small. ∎
Theorem 3.3 demonstrates that the KL divergence, a specific instance of an $f$-divergence, is not geodesically convex. This implies that, in general, $f$-divergences do not necessarily attain such convexity. In what follows, we establish a sufficient condition on $f$ that guarantees the geodesic convexity of the corresponding $f$-divergence.
Theorem 3.4.
Assume $f\in C^{2}(0,+\infty)$ is convex, satisfying $f(1)=0$ and $xf^{\prime}(x)$ is concave. Then, for any smooth, positive target density $\rho^{*}$, the $f$-divergence is geodesically convex within the Fisher-Rao geometry. Furthermore, when $f^{\prime\prime}(1)>0$, such an $f$-divergence is $\alpha_{f}$-strongly geodesically convex, where $\alpha_{f}>0$ is a constant that depends solely on $f$. These assumptions are satisfied by $f^{\prime\prime}(x)=x^{p}$ for $p\leq-2$.
Proof.
When $xf^{\prime}(x)$ is concave, using Jensen’s inequality we get
(3.25) $\mathbb{E}_{\rho}[f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}]=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\frac{\rho}{{\rho^{*}}}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\leq\mathbb{E}_{{\rho^{*}}}\bigl{[}\frac{\rho}{{\rho^{*}}}\bigr{]}f^{\prime}\Bigl{(}\mathbb{E}_{{\rho^{*}}}\bigl{[}\frac{\rho}{{\rho^{*}}}\bigr{]}\Bigr{)}=f^{\prime}(1).$
We recall the expression of Hessian operator from (3.17b)
[…]
where in the last line we used (3.25). We introduce the function
(3.27) $h(x)=2xf^{\prime\prime}(x)+f^{\prime}(x)-f^{\prime}(1)=2(xf^{\prime}(x))^{\prime}-f^{\prime}(x)-f^{\prime}(1).$
With the function $h(x)$, we can rewrite (LABEL:e:gHess) as
(3.28) $g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)\geq\frac{1}{2}\int\rho(\psi-\mathbb{E}_{\rho}[\psi])^{2}h\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
Due to our assumptions, $xf^{\prime}(x)$ and $-f(x)$ are both concave, so $(xf^{\prime}(x))^{\prime}$ and $-f^{\prime}(x)$ are non-increasing functions of $x$. It follows that $h(x)$ (see (3.27)) is non-increasing. Therefore, when $x\leq 1$, we have
$h(x)\geq h(1)=2f^{\prime\prime}(1)\geq 0,$
and when $x\geq 1$, we have
$h(x)\geq f^{\prime}(x)-f^{\prime}(1)\geq 0.$
Therefore, $h(x)\geq 0$. This together with (3.28) implies that
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)\geq 0,$
for any tangent vector $\sigma$. Thus, the $f$-divergence is geodesically convex.
Next we prove that when $f^{\prime\prime}(1)>0$, the $f$-divergence is $\alpha_{f}$-strongly geodesically convex. To do so, we need a stronger estimate on the function $h(x)$. When $f^{\prime\prime}(1)>0$, there exists $\delta_{f}>0$, such that for $x\in[1,1+\delta_{f}]$,
$\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x).$
This estimate, together with the discussions above, implies that
(3.29) $h(x)\geq\begin{cases}2f^{\prime\prime}(1)&0<x\leq 1\\ 2xf^{\prime\prime}(x)\geq f^{\prime\prime}(1)&1<x\leq 1+\delta_{f}\\ f^{\prime}(x)-f^{\prime}(1)=\int_{1}^{x}f^{\prime\prime}(\xi){\rm d}\xi\geq\delta_{f}\frac{f^{\prime\prime}(1)}{2}&x\geq 1+\delta_{f}\end{cases}.$
Therefore, $h(x)\geq\alpha_{f}>0$, where $\alpha_{f}=\min\{f^{\prime\prime}(1),\delta_{f}\frac{f^{\prime\prime}(1)}{2}\}$ is a constant that depends only on $f$. Therefore, when $f^{\prime\prime}(1)>0$, combining (3.29) and (3.28), we conclude that the $f$-divergence is $\alpha_{f}$-strongly geodesically convex, i.e.,
$g_{\rho}^{\rm FR}(\textrm{Hess}_{\rm FR}D_{f}[\rho\|{\rho^{*}}]\sigma,\sigma)\geq\frac{\alpha_{f}}{2}g_{\rho}^{\rm FR}(\sigma,\sigma).$
Finally, for $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$, we have $f$ is convex on $(0,+\infty)$. Moreover, $xf^{\prime}(x)$ is concave, since $(xf^{\prime}(x))^{\prime\prime}=2f^{\prime\prime}(x)+xf^{\prime\prime\prime}(x)=(2+p)x^{p}\leq 0$ for $x\in(0,+\infty)$. Thus, all the assumptions are satisfied for $f$ with $f^{\prime\prime}(x)=x^{p}$ with $p\leq-2$. ∎
4. Functional Inequality: Gradient Dominance Condition
In this section, we turn to the gradient dominance condition (1.11) of the $f$-divergences in the Fisher-Rao geometry. Counterexamples will be presented to illustrate that the KL-divergence does not satisfy the gradient dominance condition. Then sufficient and necessary conditions will be provided to identify the $f$ that ensures the gradient dominance condition.
First, we calculate the norm of the gradient in the Fisher-Rao geometry
$g_{\rho}^{\mathrm{FR}}(\textrm{grad}_{\rm FR}D_{f}[\rho\|{\rho^{*}}],\textrm{grad}_{\rm FR}D_{f}[\rho\|{\rho^{*}}])$
$=$ $g_{\rho}^{\mathrm{FR}}\left(-\rho f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}+\rho\mathbb{E}_{\rho}\Bigl{[}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]},-\rho f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}+\rho\mathbb{E}_{\rho}\Bigl{[}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\right)$
$=$ $\int\rho\left(f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}\Bigl{[}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\right)^{2}\mathrm{d}\theta.$
Thus, the gradient dominance condition can be formulated as
Condition 4.1 (Gradient dominance condition).
There exists $\alpha_{f}>0$ depending only on $f$, such that
(4.1) $\int\rho\left(f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-\mathbb{E}_{\rho}\Bigl{[}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}\Bigr{]}\right)^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho\|{\rho^{*}}].$
The above inequality holds for any smooth positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
4.1. KL divergence: no gradient dominance
We first present the following negative result concerning the gradient dominance condition for the KL-divergence.
Theorem 4.1.
Given any smooth, positive target distribution ${\rho^{*}}$ and any $\epsilon>0$, we can construct $\rho$, such that $\rho\neq{\rho^{*}}$ and
(4.2) $\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr{)}^{2}\mathrm{d}\theta\leq\epsilon\rm{KL}[\rho\|{\rho^{*}}].$
Moreover, given any $\epsilon^{\prime}>0$, we can also construct $\rho$ in the neighborhood of ${\rho^{*}}$ defined as KL$[\rho\|{\rho^{*}}]<\epsilon^{\prime}$, such that (4.2) holds.
Proof.
We denote
(4.3) $\mathcal{G}(\rho,{\rho^{*}})=\int\rho\Bigl{(}\log\frac{\rho}{{\rho^{*}}}-{\rm KL}[\rho\|{\rho^{*}}]\Bigr{)}^{2}\mathrm{d}\theta.$
We first consider 1D Gaussian densities, $\rho(\theta)=\mathcal{N}(\theta;\mu,\sigma)$ and ${\rho^{*}}(\theta)=\mathcal{N}(\theta;0,1)$. We can calculate both sides of (4.2):
$\mathcal{G}(\rho,{\rho^{*}})=\frac{1}{2}\Bigl{(}\sigma^{2}-1\Bigr{)}^{2}+\sigma^{2}\mu^{2},\qquad{\rm KL}[\rho\|{\rho^{*}}]=\frac{\sigma^{2}}{2}+\frac{\mu^{2}}{2}-\log\sigma-\frac{1}{2}.$
We choose $M>\max\{1,\sqrt{\frac{3}{\epsilon}}\}$ and set $\sigma=\frac{1}{M}$ and $\mu=M$, then
$\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=\frac{\frac{1}{2}\Bigl{(}\frac{1}{M^{2}}-1\Bigr{)}^{2}+1}{\frac{1}{2M^{2}}+\log M-\frac{1}{2}+\frac{M^{2}}{2}}\leq\frac{3}{M^{2}}\leq\epsilon.$
Here we used the fact that $\frac{1}{2M^{2}}+\log M\geq\frac{1}{2}$ when $M\geq 1$.
Then we consider the general case. Given any ${\rho^{*}}$, we construct
(4.4) $\rho(\theta)={\rho^{*}}(\theta)(x_{1}\mathbbm{1}_{S^{c}_{r}}(\theta)+x_{2}\mathbbm{1}_{S_{r}}(\theta))$
where $\mathbbm{1}_{A}$ is the indicator function of a set $A$, and $S_{r}$ denotes the sphere with radius $r$. Then, similar to the proof of Theorem 3.3, we have the flexibility to select any $x_{1}>1>x_{2}>0$; in particular, we can adjust $r$ to ensure $\rho$ is a density function.
We calculate both sides of eq. 4.2 as follows:
$\mathcal{G}(\rho,{\rho^{*}})=x_{1}x_{2}\frac{(x_{1}-1)(1-x_{2})}{(x_{1}-x_{2})^{2}}\Bigl{(}\log\frac{x_{1}}{x_{2}}\Bigr{)}^{2}$
${\rm KL}[\rho\|{\rho^{*}}]=\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}+\frac{x_{1}x_{2}-x_{2}}{x_{1}-x_{2}}\log x_{2}.$
[…]
(4.5a) ${\rm KL}[\rho\|{\rho^{*}}]\leq$ $\frac{x_{1}-x_{1}x_{2}}{x_{1}-x_{2}}\log x_{1}\leq\log x_{1}=\epsilon^{\prime}$
(4.5b) $\frac{\mathcal{G}(\rho,{\rho^{*}})}{{\rm KL}[\rho\|{\rho^{*}}]}=$ $\frac{(1-\frac{1}{x_{1}})(\frac{1}{x_{2}}-1)\bigl{(}\log\frac{x_{1}}{x_{2}}\bigr{)}^{2}}{(\frac{1}{x_{1}}-\frac{1}{x_{2}})\bigl{(}(\frac{1}{x_{2}}-1)\log x_{1}+(1-\frac{1}{x_{1}})\log x_{2}\bigr{)}}$
$=$ $\frac{(1-e^{-\epsilon^{\prime}})(e^{M}-1)(M+\epsilon^{\prime})^{2}}{(e^{M}-e^{-\epsilon^{\prime}})\bigl{(}(e^{M}-1)\epsilon^{\prime}-(1-e^{-\epsilon^{\prime}})M\bigr{)}}$
$<$ $\frac{(M+\epsilon^{\prime})^{2}}{(e^{M}-1)\epsilon^{\prime}-(1-e^{-\epsilon^{\prime}})M}<\frac{(2M)^{2}}{\epsilon^{\prime}e^{M}-2M}\epsilon.$
In the last inequality, we used that when $M\geq 20$, we have $e^{M/2}\geq 6M^{2}$, which implies
$4M^{2}\leq\epsilon\epsilon^{\prime}e^{M/2}6M^{2}-2M^{2}\leq(\epsilon e^{M}-2M)\epsilon^{\prime}.$
Finally, using the same mollifying approach in the proof of Theorem 3.3, we can smooth $\rho$ in (4.4) and obtain a anologous result. The proof is complete.
4.2. Sufficient and necessary conditions for gradient dominance
Moreover, we have the following sufficient and necessary condition on $f$ to satisfy the gradient dominance in Theorem 4.2. Since that both sides of the gradient dominance inequality (4.1) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$, we can further assume $f^{\prime}(1)=0$.
Condition 4.2.
There exists an $\alpha>0$, such that
(4.6) $\begin{split}\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}\end{split}$
holds for any $0<x<1<y$.
Theorem 4.2.
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$, then Condition 4.2 is sufficient and necessary for gradient dominance (4.1) to hold.
Sketch of Proof.
The proof breaks into the following three steps. In the first step, we use measure theoretical tools to reduce the gradient dominance condition to an inequality on the discrete probability simplex, which is simpler to deal with. In the second step, we use novel linear programming techniques to reduce the global nonlocal term in the inequality to a three-point interaction term. Finally, we show the two-point version of inequality, which is the sufficient and necessary condition, implies the three-point ones, thus completing the proof. We present a sketch of main techniques here and refer to the Appendix for detailed proofs of the arguments.
Step 1: Reducing to an inequality on the discrete probability simplex
We need the following lemma, whose proof is in Appendix A.
Lemma 4.3.
There exists a sequence of bounded domains $\Omega_{n}\subset\mathbb{R}^{d}$ for $n\geq 1$ such that
(4.7) $\Omega_{n}\subset\Omega_{n+1},\quad\bigcup_{n\geq 1}\Omega_{n}=\mathbb{R}^{d},\quad\text{and}\quad\int_{\Omega_{n}}\rho(\theta){\rm d}\theta=\int_{\Omega_{n}}\rho^{*}(\theta){\rm d}\theta.$
Based on Lemma 4.3, consider the truncated and normalized densities
(4.8) $\rho_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho(\theta^{\prime}){\rm d}\theta^{\prime}}\rho(\theta)\mathbbm{1}_{\Omega_{n}}(\theta)\quad\text{and}\quad\rho^{*}_{n}(\theta)=\frac{1}{\int_{\Omega_{n}}\rho^{*}(\theta^{\prime}){\rm d}\theta^{\prime}}\rho^{*}(\theta)\mathbbm{1}_{\Omega_{n}}(\theta).$
Claim: it suffices to show for all $n\geq 1$, the following holds
(4.9) $\int\rho_{n}\bigl{(}f^{\prime}\bigl{(}\frac{\rho_{n}}{\rho^{*}_{n}}\bigr{)}-c\bigr{)}^{2}\mathrm{d}\theta\geq\alpha_{f}D_{f}[\rho_{n}\|\rho^{*}_{n}],$
for any constant $c\in\mathbb{R}$. The rationale behind this claim is that, once the above hold, we can take $n\to\infty$. Here we have
(4.10) $\lim_{n\to\infty}\int\rho_{n}\big{(}f^{\prime}\bigl{(}\frac{\rho_{n}}{\rho^{*}_{n}}\bigr{)}-c\big{)}^{2}\mathrm{d}\theta=\frac{\lim_{n\to\infty}\int_{\Omega_{n}}\rho\big{(}f^{\prime}\bigl{(}\frac{\rho}{\rho^{*}}\bigr{)}-c\big{)}^{2}\mathrm{d}\theta}{\lim_{n\to\infty}\int_{\Omega_{n}}\rho(\theta){\rm d}\theta}=\int\rho\bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-c\bigr{)}^{2},$
where the convergences of the numerator and denominator are due to the monotone convergence theorem. Moreover, by Fatou’s lemma, we have $\underline{\lim}_{n\to\infty}D_{f}[\rho_{n}\|\rho^{*}_{n}]\geq D_{f}[\rho\|\rho^{*}]$. Combining the two limits together, we get
(4.11) $\int\rho\bigl{(}f^{\prime}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}-c\bigr{)}^{2}\geq\alpha D_{f}[\rho\|\rho^{*}].$
By optimizing the constant $c$, we arrive at the desired gradient dominance condition.
Therefore, it suffices to prove the claim in (4.9), which is the same inequality as (4.1) but applied to densities in bounded domains. Suppose $\Omega_{n}=\cup_{k=1}^{K}\omega_{k}$ is partitioned to $K$ disjoint subdomains. We can approximate $\rho_{n}$ and $\rho^{*}$ by
$\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}\quad\text{and}\quad\sum_{k=1}^{K}\frac{\int_{\omega_{k}}\rho_{n}^{*}{\rm d}\theta}{\int_{\omega_{k}}{\rm d}\theta}\mathbbm{1}_{\omega_{k}}.$
If we can prove the inequality for the above approximate densities, we can then take finer and finer partitions and as $K\to\infty$, we will get the corresponding inequality for $\rho_{n}$ and $\rho^{*}$. This is true because we work in a bounded domain $\Omega_{n}$, for which we can apply uniform convergence to establish the limit; note that $\rho$ and $\rho^{*}$ are assumed smooth.
As a consequence, it suffices to prove the inequality when the densities are summation of indicator functions. By an abuse of notations, we denote the corresponding mass of the above approximations in each support of the indicator functions by $\rho_{i},\rho_{i}^{*}$ for $1\leq i\leq K$. Furthermore, denote by $x_{i}=\frac{\rho_{i}}{\rho_{i}^{*}}$. It suffices to show
(4.12) $\sum_{i=1}^{K}\rho_{i}\bigl{(}f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr{)}^{2}\geq\alpha_{f}\sum_{i}\frac{\rho_{i}}{x_{i}}f(x_{i}),$
for any $\{\rho_{i}\}_{i=1}^{K}$ and $\{x_{i}\}_{i=1}^{K}$ that satisfy
(4.13) $\sum_{i=1}^{K}\rho_{i}=1,\quad\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
This can be viewed as a reduced gradient dominance condition for probability densities on a discrete simplex.
Step 2: Reducing the nonlocal term to a three-point interaction
First, we derive an equivalent form of the inequality. For the left hand side, we have
(4.14) $\sum_{i=1}^{K}\rho_{i}\bigl{(}f^{\prime}(x_{i})-\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j})\bigr{)}^{2}=\sum_{i=1}^{K}\rho_{i}(f^{\prime}(x_{i}))^{2}-(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j}))^{2}$
$=$ $\frac{1}{2}\left(\sum_{i=1}^{K}\rho_{i}(f^{\prime}(x_{i}))^{2}+\sum_{i=j}^{K}\rho_{j}(f^{\prime}(x_{j}))^{2}-2(\sum_{i=1}^{K}\rho_{j}f^{\prime}(x_{i}))(\sum_{j=1}^{K}\rho_{j}f^{\prime}(x_{j}))\right)$
$=$ $\frac{1}{2}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}\bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\bigr{)}^{2}.$
For the right hand side, we get
(4.15) $\begin{split}\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}f(x_{i})&=\sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{i}\frac{\rho_{i}}{1-x_{i}}f(x_{i})\\ &=\sum_{1\leq i,j\leq K}\frac{\rho_{i}\rho_{j}}{x_{i}(1-x_{i})}f(x_{i})-\sum_{ij}\frac{\rho_{i}\rho_{j}}{x_{j}(1-x_{i})}f(x_{i})\\ &=\frac{1}{2}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{j}}-\frac{1}{x_{i}})\bigl{(}\frac{f(x_{j})}{1-x_{j}}-\frac{f(x_{i})}{1-x_{i}}\bigr{)},\end{split}$
where to get the second line we used (4.13). We remark that since $f\in C^{2}(0,+\infty)$ and $f(1)=0$, $f(x)/(1-x)$ is well defined and there is no singularity at $x=1$. Therefore, the equivalent inequality is
(4.16) $\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq K}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
When $K=2$, denote $x=x_{1}$ and $y=x_{2}$ with $0<x<1<y$. We have $\rho_{1}=\frac{x(y-1)}{y-x}$ and $\rho_{2}=\frac{y(1-x)}{y-x}$, (4.16) leads to
(4.17) $\begin{split}\exists\ \alpha>0\textrm{ s.t. }&\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ 0<x<1<y\end{split}$
with $\alpha=\alpha_{f}$. This is the same as Condition 4.2. It is a necessary condition for the discrete inequality, and is also necessary for the continuous inequality by smoothing the indicator functions. It remains to show that Condition 4.2 is also sufficient, i.e., the case $K=2$ holds implies that the inequality holds for any $K\geq 3$.
We first show that cases $K\geq 3$ can be reduced to $K=3$. To do so, we use the original formulation (4.12). We fix $x_{i}>0$, $\alpha_{f}>0$, and $M=\sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})$, and consider the following linear programming problem with respect to variables $\{\rho_{i}\}_{i=1}^{K}$
$\min$ $\sum_{i=1}^{K}\rho_{i}\Bigl{(}f^{\prime}(x_{i})^{2}-\alpha_{f}\frac{f(x_{i})}{x_{i}}\Bigr{)}-M^{2}$
s.t. $\sum_{i=1}^{K}\rho_{i}=1,\ \sum_{i=1}^{K}\frac{\rho_{i}}{x_{i}}=1,\ \sum_{i=1}^{K}\rho_{i}f^{\prime}(x_{i})=M,\ \rho_{i}\geq 0.$
The linear programming problem has a feasible solution and the constraint set is bounded. We claim that there is a minimizer that has at most $3$ nonzero $\rho_{i}$. Otherwise, without loss of generality, we assume for the minimizer $\rho$, it holds $\rho_{j}>0$ for $1\leq j\leq 4$. We can adjust them in the direction $\vec{d}=[d_{1},\,d_{2},\,d_{3},\,d_{4}]$ satisfying $\sum_{j=1}^{K}d_{j}=0$, $\sum_{j=1}^{K}\frac{d_{j}}{x_{j}}=0$, and $\sum_{j=1}^{K}d_{j}f^{\prime}(x_{j})=0$ to reduce the objective function or increase the number of nonzero $\rho_{i}$. This leads to a contradition so the claim is true. This means that we only need to prove the inequality for the case $K=3$. Using the equivalent formulation (4.16), assuming (4.6) holds, we need to show
(4.18) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
Step 3: Showing the three-point inequality is implied by two-point ones
The last step is to use the inequality in the case of $K=2$ to prove (4.18). We have the following two technical lemmas whose proofs are in Appendix A.
Lemma 4.4.
Assume (4.6) holds. There exists $10<\delta<$ 1, and $0<\alpha_{\delta}\leq\alpha$, such that
(4.19) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ x\in(1-\delta,1+\delta)\ \textrm{and}\ y>0.$
Lemma 4.5.
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime}(1)=0$ and under the constraint (4.13), conditions (4.6) and (4.19) implies (4.18).
With the above lemmas, the sufficient condition part is proved. ∎
4.3. An easier-to-verify sufficient condition
We provide a simpler and easier-to-verify sufficient condition on $f$ to ensure the gradient dominance 4.1 holds.
Theorem 4.6.
We assume $f\in C^{2}(0,+\infty)$ is a convex function with $f(1)=0$ and there exists $\alpha_{s}>0$ such that $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$. Then Condition 4.2 holds. Note that the assumptions are satisfied with $f^{\prime\prime}(x)=x^{p}$ and $p\leq-2$.
Proof.
The assumption $x^{2}f^{\prime\prime}(x)>\alpha_{s}$ for $0<x\leq 1$ implies that $f^{\prime\prime}(1)>\alpha_{s}$. Combining this with the condition $f\in C^{2}$ leads to the existence of $0<\delta<\frac{1}{2}$, such that
(4.20) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\quad\text{for}\quad 1-2\delta\leq x\leq 1+2\delta.$
The convexity of $f(x)$ leads to
(4.21) $(x-1)f^{\prime}(x)\geq f(x)-f(1)=f(x),\qquad\forall x>0.$
With the above preparations, we prove that for $0<x<1<y$, Condition 4.2 holds with $\alpha=\alpha_{s}\delta$. This can be obtained from the following two inequalities.
[…]
$\geq\alpha_{s}\bigl{(}\frac{1}{x}-1\bigr{)}+\delta f^{\prime\prime}(1)\bigl{(}1-\frac{1}{y}\bigr{)}$
$\geq\alpha_{s}\delta\bigl{(}\frac{1}{x}-\frac{1}{y}\bigr{)}\quad(\textrm{Using }f^{\prime\prime}(1)\geq\alpha_{s})$
[…]
The proof is complete. ∎
Remark 4.7.
Recall that in general, the gradient dominance condition can be implied by strong geodesic convexity of the energy functional. In Theorem 3.4, a sufficient condition for strong geodesic convexity is that $xf^{\prime}(x)$ is concave and $f^{\prime\prime}(1)>0$. If $xf^{\prime}(x)$ is concave, we can show that $x^{2}f^{\prime\prime}(x)\geq f^{\prime\prime}(1)$ for any $0<x\leq 1$. To do so, we rewrite $x^{2}f^{\prime\prime}(x)$ in terms of $xf^{\prime}(x)$ as $x^{2}f^{\prime\prime}(x)=x\bigl{(}xf^{\prime}(x)\bigr{)}^{\prime}-xf^{\prime}(x).$ If $xf^{\prime}(x)$ is concave, then $h(x):=\bigl{(}xf^{\prime}(x)\bigr{)}^{\prime}=xf^{\prime\prime}(x)+f^{\prime}(x)$ is nonincreasing. We have
$x^{2}f^{\prime\prime}(x)-f^{\prime\prime}(1)$ $=\bigl{(}xh(x)-xf^{\prime}(x)\bigr{)}-\bigl{(}h(1)-f^{\prime}(1)\bigr{)}$
$=\Bigl{(}\int_{0}^{x}h(x){\rm d}\xi-\int_{0}^{x}h(\xi)\mathrm{d}\xi\Bigr{)}-\Bigl{(}\int_{0}^{1}h(1){\rm d}\xi-\int_{0}^{1}h(\xi)\mathrm{d}\xi\Bigr{)}$
$=\int_{0}^{x}h(x){\rm d}\xi+\int_{x}^{1}h(\xi)\mathrm{d}\xi-\int_{0}^{1}h(1){\rm d}\xi$
$=\int_{0}^{x}h(x)-h(1)\mathrm{d}\xi+\int_{x}^{1}h(\xi)-h(1)\mathrm{d}\xi$
$\geq 0\quad(\textrm{Using }h(1)\leq h(x),h(\xi)).$
Therefore, Theorem 4.6 is satisfied with $\alpha_{s}=f^{\prime\prime}(1)$, so the gradient dominance holds. This establishes a connection between Theorem 4.6 and Theorem 3.4.
5. Functional Inequality: Dual Gradient Dominance Condition
The KL divergence is the most widely used energy functional for sampling; however, the negative results in Theorems 3.3 and 4.1 imply that the concepts of geodesic convexity and the gradient dominance condition are not sufficient for analyzing the convergence of the Fisher-Rao gradient flow of the KL divergence.
In this section, we introduce a dual gradient dominance condition, providing a more powerful functional inequality to understand the convergence behavior of the Fisher-Rao gradient flow associated with various $f$-divergences. To the best of our knowledge, such dual functional inequalities have not been explored in the Wasserstein setting, making this a novel and non-trivial discovery in the context of the Fisher-Rao geometry.
5.1. Non-increasing of general $f$-divergences
Take a general convex function ${\bar{f}}\in C^{2}(0,+\infty),{\bar{f}}(1)=0$, we consider the following $\bar{f}$-divergence
(5.1) $D_{\bar{f}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta.$
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, we have that
(5.2) $\frac{\mathrm{d}}{\mathrm{d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]=\int\partial_{t}\rho_{t}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
$=$ $\int\Bigl{(}-\rho_{t}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})+\rho_{t}\mathbb{E}_{\rho_{t}}\bigl{[}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})\bigr{]}\Bigr{)}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}}{{\rho^{*}}}\bigr{)}\mathrm{d}\theta$
$=$ $\int\Bigl{(}-\rho_{t}(\theta)f^{\prime}(\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)})+\rho_{t}(\theta)\int\rho_{t}(\theta^{\prime})f^{\prime}(\frac{\rho_{t}(\theta^{\prime})}{{\rho^{*}}(\theta^{\prime})})\mathrm{d}\theta^{\prime}\Bigr{)}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)}\bigr{)}\mathrm{d}\theta$
$=$ $\int\int\rho_{t}(\theta)\rho_{t}(\theta^{\prime})\Bigl{(}-f^{\prime}(\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)})+f^{\prime}(\frac{\rho_{t}(\theta^{\prime})}{{\rho^{*}}(\theta^{\prime})})\Bigr{)}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)}\bigr{)}\mathrm{d}\theta^{\prime}\mathrm{d}\theta$
$=$ $\frac{1}{2}\int\int\rho_{t}(\theta)\rho_{t}(\theta^{\prime})\Bigl{(}-f^{\prime}(\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)})+f^{\prime}(\frac{\rho_{t}(\theta^{\prime})}{{\rho^{*}}(\theta^{\prime})})\Bigr{)}\Bigl{(}{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}(\theta)}{{\rho^{*}}(\theta)}\bigr{)}-{\bar{f}}^{\prime}\bigl{(}\frac{\rho_{t}(\theta^{\prime})}{{\rho^{*}}(\theta^{\prime})}\bigr{)}\Bigr{)}\mathrm{d}\theta^{\prime}\mathrm{d}\theta$
$\leq$ $0$
This shows that any $D_{\bar{f}}$ is non-increasing along the Fisher-Rao gradient flow with respect to $D_{f}$. Here in the last inequality, we used the fact that both $f^{\prime}(x)$ and ${\bar{f}}^{\prime}(x)$ are non-decreasing functions.
5.2. Dual gradient dominance condition
Motivated by the non-increasing of $D_{\bar{f}}$ for any ${\bar{f}}$, in the following we introduce the dual gradient dominance condition. The idea is that we require the derivative in (5.2) to be lower bounded by a constant multiplying the sum of $D_{f}$ and $D_{\bar{f}}$, similar to the gradient dominance condition but applied to a different energy functional $D_{\bar{f}}$.
Condition 5.1 (Dual gradient dominance condition).
There exists $\alpha_{f}>0$ depending only on $f$, such that
(5.3) $\begin{split}-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}{\bar{f}}^{\prime}(\frac{\rho}{{\rho^{*}}})\mathrm{d}\theta\geq\alpha_{f}\Bigl{(}D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr{)}.\end{split}$
The above inequality holds for any positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
Remark 5.1.
If the dual gradient dominance condition holds along the gradient flow of $D_{f}$, we will get
$\frac{{\rm d}}{{\rm d}t}(D_{f}[\rho_{t}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{t}\|{\rho^{*}}])\leq-\alpha_{f}(D_{f}[\rho_{t}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]),$
which implies the exponential convergence, i.e., $\mathcal{O}(\exp(-\alpha_{f}t))$, of $D_{f}[\rho_{t}\|{\rho^{*}}]+D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]$. $\lozenge$
In the following, we will consider two special choices of ${\bar{f}}$:
  1. (1)
    ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$ (corresponding to the reverse $\chi^{2}$ divergence)
  2. (2)
    ${\bar{f}}(x)=xf(\frac{1}{x})$ (corresponding to the *-conjugate of $f$ [67]).
We prove the dual gradient dominance condition holds under mild assumptions on $f$. Specifically, the first choice leads to a general result, while the second choice leads to tighter results for specific $f$ including the KL divergence.
5.3. General results using the reverse $\chi^{2}$ divergence as the dual
We take ${\bar{f}}(x)=\frac{(x-1)^{2}}{x}$, the $f$-divergence associated with ${\bar{f}}$ is the reverse $\chi^{2}$ distance
$D_{\bar{f}}[\rho\|{\rho^{*}}]=\chi^{2}[{\rho^{*}}\|\rho]=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta-1.$
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, the derivative (5.2) is given explicitly as
$\frac{\mathrm{d}}{\mathrm{d}t}\chi^{2}[{\rho^{*}}\|\rho_{t}]$ $=\int\partial_{t}\rho_{t}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho_{t}}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta$
$=\int\Bigl{(}-\rho_{t}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})+\rho_{t}\mathbb{E}_{\rho_{t}}\bigl{[}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho_{t}}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta.$
For this choice of ${\bar{f}}=\frac{(x-1)^{2}}{x}$, the generalized gradient dominance condition holds, which follows from the following theorem.
Theorem 5.2.
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime\prime}(1)>0$, there exists an $\alpha_{f}>0$ that depends on $f$ only, such that
(5.4a) $-\int\Bigl{(}-\rho f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\bigr{]}\Bigr{)}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta\geq D_{f}[\rho\|{\rho^{*}}],$
(5.4b) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}1-\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{2}\Bigr{)}\mathrm{d}\theta\geq 2\alpha_{f}\chi^{2}[{\rho^{*}}\|\rho].$
The above inequalities hold for any positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
Remark 5.3.
As a consequence of Theorem 5.2, the Fisher-Rao gradient flow (2.3) associated with any $f$-divergence, assumed well-posed, converges exponentially fast such that
$D_{f}[\rho_{t}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{t}]\leq e^{-\alpha_{f}t}\Bigl{(}D_{f}[\rho_{0}\|{\rho^{*}}]+\chi^{2}[{\rho^{*}}\|\rho_{0}]\Bigr{)}.$
The investigation of the well-posedness of the equation is left as a future work. $\lozenge$
Proof for Theorem 5.2.
We denote
$\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\geq 1,$
then, the left hand side of (5.4a) or (5.4b) can be rewritten as
(5.5) $\begin{split}-\int\Bigl{(}-\rho f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}&+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\bigr{]}\Bigr{)}\Bigl{(}1-\frac{{\rho^{*}}^{2}}{\rho^{2}}\Bigr{)}\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta+\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta\cdot\int{\rho}f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta\\ =&-\int f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ =&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta.\end{split}$
To proceed, we rely in the following local inequality. Its proof is in Appendix B.
Lemma 5.4.
Denote $\mu^{2}=\int\frac{{\rho^{*}}^{2}}{\rho}\mathrm{d}\theta$, $\mu>0$, then for convex $f$ and $\frac{\rho^{*}}{\rho}\in(0,+\infty)$, it holds that
(5.6) $\Bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})).$
Multiplying the local inequality (5.6) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
$\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}$ $\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta$
(5.7) $\geq$ $\int\Bigl{(}\mu({\rho^{*}}-\mu\rho)f^{\prime}(\frac{1}{\mu})-\mu^{2}{\rho^{*}}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}}))\Bigr{)}\mathrm{d}\theta$
$=$ $\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta$
$\geq$ $D_{f}[\rho\|{\rho^{*}}].$
In the last inequality, we used the fact that $0=f(1)\geq f(\frac{1}{\mu})-(\frac{1}{\mu}-1)f^{\prime}(\frac{1}{\mu})$ by the convexity of $f$; note that $\mu\geq 1$. This completes the proof of (5.4a).
To show (5.4b), we first introduce another local inequality as follows; its proof is in Appendix B.
Lemma 5.5.
Given a convex function $f\in C^{2}(0,+\infty)$ with $f(1)=0$ and $f^{\prime\prime}(1)>0$, there exists $0<\delta_{f}<\frac{1}{4}$, such that
(5.8) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x),\quad\text{for any }x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}}).$
Moreover, there exists $\alpha_{f}^{\prime}>0$, such that
(5.9) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}(x-\mu)^{2}$
for any $x>0$ and $\mu\in[1,1+\delta_{f}]$.
We will employ the quantities $\delta_{f}$ and $\alpha^{\prime}_{f}$ defined in Lemma 5.5. We prove the dual gradient dominance condition (5.4b) for the following two scenarios, depending on whether $\mu$ is greater than $1+\delta_{f}$ or not.
When $\mu\leq 1+\delta_{f}$, taking $x={\rho^{*}}/\rho$ in Lemma 5.5 leads to
(5.10) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{)}(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2})\geq\alpha_{f}^{\prime}(\frac{{\rho^{*}}}{\rho}-\mu)^{2}.$
Multiplying the local inequality (5.10) by $\rho$ and integrating over $\mathbb{R}^{d}$ leads to
(5.11) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\geq\alpha^{\prime}_{f}\int(\frac{{\rho^{*}}}{\rho}-\mu)^{2}\rho\mathrm{d}\theta\\ =&\alpha^{\prime}_{f}\int\frac{{\rho^{*}}^{2}}{\rho}+\rho\mu^{2}-2{\rho^{*}}\mu\mathrm{d}\theta=2\alpha^{\prime}_{f}\mu(\mu-1)\quad(\textrm{using }\mu\geq 1)\\ \geq&\alpha^{\prime}_{f}(\mu+1)(\mu-1)=\alpha^{\prime}_{f}\chi^{2}[{\rho^{*}}\|\rho].\\ \end{split}$
When $\mu\geq 1+\delta_{f}$, utilizing (5.7) yields
(5.12) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})+\mu^{2}\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu}).\end{split}$
where in the last inequality, we used that the fact that the $f$-divergence is non-negative. Let denote $h(x)=(x-1)f^{\prime}(x)-f(x)$. Then for $x\in(0,1]$, $h(x)$ is nonnegative and non-increasing, since $h^{\prime}(x)=(x-1)f^{\prime\prime}(x)\leq 0$ and $h(1)=0$. Therefore, the inequality (LABEL:eq:generalized-gradient-dominant-condition-2b0) can be further carried on to obtain
(5.13) $\begin{split}&\int\Bigl{(}f^{\prime}\Bigl{(}\frac{1}{\mu}\Bigr{)}-f^{\prime}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\rho\mathrm{d}\theta\\ \geq&\mu(1-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}f(\frac{1}{\mu})=\mu^{2}h(\frac{1}{\mu})\quad(\textrm{using }\mu\geq 1+\delta_{f}\textrm{ and }h\textrm{non-increasing},h\geq 0)\\ \geq&h(\frac{1}{1+\delta_{f}})(\mu^{2}-1)=h(\frac{1}{1+\delta_{f}})\chi^{2}[{\rho^{*}}\|\rho].\end{split}$
According to the definition of $\delta_{f}$ in (5.9), we have
(5.14) $h(\frac{1}{1+\delta_{f}})=\int^{1}_{\frac{1}{1+\delta_{f}}}f^{\prime\prime}(x)(1-x)\mathrm{d}x\geq\frac{f^{\prime\prime}(1)}{2}\int^{1}_{\frac{1}{1+\delta_{f}}}(1-x)\mathrm{d}x=\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}.$
Combining the two scenarios, namely $\mu\leq 1+\delta_{f}$ in (5.11) and $\mu\geq 1+\delta_{f}$ in (LABEL:eq:generalized-gradient-dominant-condition-2b), we conclude that (5.4b) holds for
$\alpha_{f}=\frac{1}{2}\min\{1,\alpha^{\prime}_{f},\frac{\delta_{f}^{2}f^{\prime\prime}(1)}{4(1+\delta_{f})^{2}}\}>0.$
This completes the proof. ∎
5.4. Tighter results using the *-conjugate $f$-divergence as the dual
In the last subsection, we obtain general convergence results by choosing $\bar{f}$ as the reverse $\chi^{2}$ divergence. The constant $\alpha_{f}$ depends on $f$ only. However, it depends on some implicit constant such as $\delta_{f},\alpha_{f}^{\prime}$ that are dependent on $f$. The goal of this subsection is to improve the estimate for specific cases and obtain more explicit constants.
We take ${\bar{f}}(x)=xf\bigl{(}\frac{1}{x}\bigr{)}$, known as as the *-conjugate function of $f$ [67]. For this choice, the $\bar{f}$-divergence has the following dual form
(5.15) $D_{{\bar{f}}}[\rho\|{\rho^{*}}]=\int{\rho^{*}}{\bar{f}}\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta=\int\rho f\Bigl{(}\frac{{\rho^{*}}}{\rho}\Bigr{)}\,\mathrm{d}\theta.$
We note that ${\bar{f}}(x)=xf\bigl{(}\frac{1}{x}\bigr{)}$ is also convex with ${\bar{f}}(1)=0$.
Along the Fisher-Rao gradient flow (2.3) with respect to $D_{f}$, we have
$\frac{\mathrm{d}}{\mathrm{d}t}D_{\bar{f}}[\rho_{t}\|{\rho^{*}}]$ $=\int\partial_{t}\rho_{t}\Bigl{(}f(\frac{{\rho^{*}}}{\rho_{t}})-\frac{{\rho^{*}}}{\rho_{t}}f^{\prime}(\frac{{\rho^{*}}}{\rho_{t}})\Bigr{)}\mathrm{d}\theta$
$=\int\Bigl{(}-\rho_{t}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})+\rho_{t}\mathbb{E}_{\rho_{t}}\bigl{[}f^{\prime}(\frac{\rho_{t}}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho_{t}})-\frac{{\rho^{*}}}{\rho_{t}}f^{\prime}(\frac{{\rho^{*}}}{\rho_{t}})\Bigr{)}\mathrm{d}\theta.$
In the following theorem, we show the dual gradient dominance condition holds with an explicit constant, when $f^{\prime\prime}$ is a polynomial.
Theorem 5.6.
Given any convex function $f$ specified by $f^{\prime\prime}(x)=x^{p}$ with $f(1)=0$, when $p\leq-2$ or $-1\leq p$, we have
(5.16) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta\geq D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}],$
and when $-2<p<-1$, we have
(5.17) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta\geq\frac{1}{2}\Bigl{(}D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr{)}.$
The above inequalities hold for any positive densities $\rho$ and ${\rho^{*}}$, provided that $f^{\prime}(\frac{\rho}{{\rho^{*}}})$ is absolutely integrable with respect to $\rho$.
Remark 5.7.
As a consequence of Theorem 5.6, we get a convergence result for the Fisher-Rao gradient flow associated with the KL-divergence. More precisely, for the Fisher-Rao gradient flow of the KL-divergence
(5.18) $\frac{\partial\rho_{t}}{\partial t}=-\rho_{t}\Bigl{(}\bigl{(}\log\rho_{t}-\log{\rho^{*}}\bigr{)}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)},$
if we choose an initial density $\rho_{0}$ satisfying that
${\rm KL}[\rho_{0}\|{\rho^{*}}]+{\rm KL}[{\rho^{*}}\|\rho_{0}]<\infty,$
then the Fisher-Rao gradient flow (assuming it is well-posed) converges exponentially fast with
${\rm KL}[\rho_{t}\|{\rho^{*}}]+{\rm KL}[{\rho^{*}}\|\rho_{t}]\leq e^{-t}\Bigl{(}{\rm KL}[\rho_{0}\|{\rho^{*}}]+{\rm KL}[{\rho^{*}}\|\rho_{0}]\Bigr{)}.$
$\lozenge$
Proof of Theorem 5.6.
First, we observe that both sides of these dual gradient dominance inequalities (5.16) and (5.17) remain the same under the transformation $f(x)\rightarrow f(x)+c(x-1)$ for any $c$. Therefore, for each $p$, it suffices to prove that (5.16) and (5.17) hold for one representative element under the transformation.
When $p=-1$, we consider $f(x)=x\log x$, the corresponding $f$-divergence is the KL-divergence. We have
(5.19) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
$=$ $\int\Bigl{(}-\rho\log(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}\log(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\frac{{\rho^{*}}}{\rho}\mathrm{d}\theta$
$=$ $\mathbb{E}_{{\rho^{*}}}\bigl{[}\log(\frac{{\rho^{*}}}{\rho})\bigr{]}+\mathbb{E}_{\rho}\bigl{[}\log(\frac{\rho}{{\rho^{*}}})\bigr{]}=D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}].$
Hence, (5.16) holds.
When $p=-2$, we consider $f(x)=-\log x$, the corresponding $f$-divergence is the reverse KL-divergence. We have
(5.20) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
$=$ $-\int\bigl{(}{\rho^{*}}-\rho\bigr{)}\bigl{(}1-\log\frac{{\rho^{*}}}{\rho}\bigr{)}\mathrm{d}\theta$
$=$ $\mathbb{E}_{{\rho^{*}}}\bigl{[}\log(\frac{{\rho^{*}}}{\rho})\bigr{]}-\mathbb{E}_{\rho}\bigl{[}\log(\frac{{\rho^{*}}}{\rho})\bigr{]}=D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}].$
Hence, (5.16) holds.
When $p\neq-1$ and $p\neq-2$, we consider $f(x)=\frac{x^{p+2}-1}{(p+2)(p+1)}$. The corresponding $f$-divergences include $\chi^{2}$ distance (when $p=0$). We have
(5.21) $-\int\Bigl{(}-\rho f^{\prime}(\frac{\rho}{{\rho^{*}}})+\rho\mathbb{E}_{\rho}\bigl{[}f^{\prime}(\frac{\rho}{{\rho^{*}}})\bigr{]}\Bigr{)}\Bigl{(}f(\frac{{\rho^{*}}}{\rho})-\frac{{\rho^{*}}}{\rho}f^{\prime}(\frac{{\rho^{*}}}{\rho})\Bigr{)}\mathrm{d}\theta$
$=$ $\frac{1}{(p+1)(p+2)}\int\Bigl{(}-\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}+\rho\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\Bigr{)}\Bigl{(}\frac{1}{p+1}+\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{p+2}\Bigr{)}\mathrm{d}\theta$
$=$ $\frac{1}{(p+2)(p+1)}\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}.$
When $p<-2$ or $-1<p$, applying Jensen’s inequality to the convex function $x^{p+2}~{}(x>0)$, we have
$\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}^{p+2}\Bigr{]}\geq 1\quad\textrm{ and }\quad\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta=\mathbb{E}_{\rho}\Bigl{[}\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{p+2}\Bigr{]}\geq 1.$
Therefore, (5.16) holds using the following estimate:
(5.22) $\frac{1}{(p+2)(p+1)}$ $\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
$\geq\frac{1}{(p+2)(p+1)}\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta-1+\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
$=D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}].$
When $-2<p<-1$, applying Jensen’s inequality to the concave function $x^{p+2}~{}(x>0)$, we have
$\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta=\mathbb{E}_{{\rho^{*}}}\Bigl{[}\bigl{(}\frac{\rho}{{\rho^{*}}}\bigr{)}^{p+2}\Bigr{]}\leq 1\quad\textrm{ and }\quad\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta=\mathbb{E}_{\rho}\Bigl{[}\bigl{(}\frac{{\rho^{*}}}{\rho}\bigr{)}^{p+2}\Bigr{]}\leq 1.$
Note that $(p+1)(p+2)<0$. We get that (5.17) using the following estimate:
(5.23) $\frac{1}{(p+2)(p+1)}$ $\Bigl{(}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
$\geq\frac{1}{(p+2)(p+1)}\Bigl{(}\frac{1}{2}\int\frac{\rho^{p+2}}{{\rho^{*}}^{p+1}}\mathrm{d}\theta+\frac{1}{2}\int\frac{{\rho^{*}}^{p+2}}{\rho^{p+1}}\mathrm{d}\theta-1\Bigr{)}$
$=\frac{1}{2}\Bigl{(}D_{\bar{f}}[\rho\|{\rho^{*}}]+D_{f}[\rho\|{\rho^{*}}]\Bigr{)}.$
This concludes the proof. ∎
6. Conclusion
In this study, we have investigated the geodesic convexity and functional inequalities related to Fisher-Rao gradient flows associated with various $f$-divergences. We have constructed counterexamples to demonstrate that, in the Fisher-Rao geometry, $f$-divergences generally lack geodesic convexity or fail to satisfy the gradient dominance condition. A notable example is the KL divergence, which is widely used for sampling applications. Nevertheless, we have also developed sufficient conditions on $f$ under which $f$-divergences possess these properties.
Furthermore, we have introduced the dual gradient dominance condition and demonstrated its validity for a broad class of $f$-divergences, notably including the KL divergence. The constants in this condition depend solely on $f$, independent of the target distribution. This characteristic sharply contrasts with the log-Sobolev inequality developed in the Wasserstein geometry, where constants depend on the target distribution.
As a direct consequence, assuming the well-posedness of the Fisher-Rao gradient flow, our results indicate that the flow exhibits exponential convergence, with a convergence rate independent of the target density.
Future research directions include establishing the well-posedness of the Fisher-Rao gradient flows and exploring the properties of more general, including non-smooth, energy functionals within the Fisher-Rao geometry and the related Wasserstein-Fisher-Rao geometry. Designing efficient numerical approximations for these Fisher-Rao gradient flows, particularly for sampling applications as discussed in [59, 60, 44, 56, 22, 63], remains a promising and open area of research.
Acknowledgments JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the EPSRC grants EP/T022132/1 and EP/V051121/1 and by the “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M, funded by MCIN/AEI/10.13039/501100011033/. YC was supported by the Courant instructorship. DZH was supported by “The Fundamental Research Funds for the Central Universities, Peking University”. We thank De Huang for useful discussions on log-Sobolev inequalities.
Appendix A Lemmas for Proof of Theorem 4.2
In this section, we prove Lemma 4.5, Lemma 4.4, and Lemma 4.3, which are needed to prove Theorem 4.2.
Proof of Lemma 4.5.
In this lemma, our assumptions (conditions (4.6) and (4.19)) are that there exists $10<\delta<$ 1, and $0<\alpha_{\delta}\leq\alpha$, such that
(A.1) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)},\ \forall\ x\in(1-\delta,1+\delta)\ \textrm{and}\ y>0,$
where the above inequality also holds for any $0<x<1<y$. We also have the constraint
(A.2) $\sum_{i=1}^{3}\rho_{i}=1,\quad\sum_{i=1}^{3}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
Our goal is to show
(A.3) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl{(}f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr{)}^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{x_{i}-1}-\frac{f(x_{j})}{x_{j}-1}\Bigr{)}.$
We assume that $x_{1}\leq x_{2}\leq x_{3}$, then due to the constraint (A.2), we have either $x_{1}=x_{2}=x_{3}=1$ or $x_{1}<1<x_{3}$.
For the first scenario, (A.3) holds with any $\alpha_{f}$. For the second scenario, when $1-\delta<x_{2}<1+\delta$, the relationships between pairs $(x_{1},x_{2})$, $(x_{1},x_{3})$ and $(x_{1},x_{3})$ fall into the conditions described in (A.1). Thus, (A.3) holds with $\alpha_{f}=\alpha_{\delta}$. Now, we only need to verify the following cases:
  1. Case 1
    . $0<x_{1}<1<1+\delta\leq x_{2}\leq x_{3}$,
  2. Case 2
    . $0<x_{1}\leq x_{2}\leq 1-\delta<1<x_{3}$.
For Case 1, we will first show that
(A.4) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}\geq\delta\rho_{2}\rho_{3}(\frac{1}{x_{2}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{2})}{1-x_{2}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}.$
In fact, the convexity of $f(x)$ implies that $\frac{f(x)}{1-x}$ is decreasing. Hence eq. A.4 can be obtained if we can show $\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\geq\delta\rho_{2}(\frac{1}{x_{2}}-\frac{1}{x_{3}}).$ This can be achieved using the following estimate:
$\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\geq\rho_{1}(\frac{1}{x_{1}}-1)=\rho_{2}(1-\frac{1}{x_{2}})+\rho_{3}(1-\frac{1}{x_{3}})\geq\rho_{2}(1-\frac{1}{x_{2}})\geq\delta\rho_{2}(\frac{1}{x_{2}}-\frac{1}{x_{3}}).$
Here in the last inequality, we used the fact that $x_{2}\geq 1+\delta$ and $x_{3}>0$.
With (A.4), we can prove that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. To do so, we denote
(A.5) $F_{i,j}=\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl{(}\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr{)},$
then we have
[…]
which shows that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$.
For Case 2, we will first show that
(A.6) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr{)}\geq\delta\rho_{2}\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{2}})\Bigl{(}\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{2})}{1-x_{2}}\Bigr{)}.$
In fact, the convexity of $f(x)$ implies that $\frac{f(x)}{1-x}$ is decreasing, and hence (A.6) can be obtained if we have $\rho_{3}\geq\delta\rho_{2}$. This can be derived by the following estimate:
$\rho_{3}\geq\rho_{3}(1-\frac{1}{x_{3}})=\rho_{1}(\frac{1}{x_{1}}-1)+\rho_{2}(\frac{1}{x_{2}}-1)\geq\rho_{2}(\frac{1}{x_{2}}-1)\geq\delta\rho_{2}.$
Here, in the last inequality, we used the fact that $x_{2}\leq 1-\delta$. Next, we will prove that (A.3) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. Using the definition in (A.5), we have
[…]
Combining both Case 1 and Case 2, we complete the proof. ∎
Proof of Lemma 4.4.
Assume that 4.2 holds with $\alpha>0$, namely
(A.7) $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}$
holds for any $0<x<1<y$. Let $y$ in (A.7) approach $1$, we get
(A.8) $f^{\prime}(x)^{2}\geq\alpha\frac{f(x)}{x}\quad\textrm{ for }\quad x>0.$
Here we used the fact that $\lim_{x\rightarrow 1}\frac{f(x)}{1-x}=\lim_{x\rightarrow 1}f^{\prime}(x)=0$. Now, letting $x$ further approach $1$ leads to that
(A.9) $f^{\prime\prime}(1)\geq\frac{\alpha}{2}.$
Here we used the fact that $\lim_{x\rightarrow 1}\frac{xf^{\prime}(x)^{2}}{f(x)}=\lim_{x\rightarrow 1}2xf^{\prime\prime}(x)+f^{\prime}(x)=2f^{\prime\prime}(1)$. Since $f^{\prime\prime}(1)>0$ and $f\in C^{2}$, there exists $0<\delta^{\prime}<\frac{1}{2}$, such that when $1-2\delta^{\prime}\leq x\leq 1+2\delta^{\prime}$,
(A.10) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\leq 2f^{\prime\prime}(1).$
Moreover, $f^{\prime}(1)=0$ implies that there exists $0<\delta<\delta^{\prime}$, such at
(A.11) $-f^{\prime}(1-\delta)\leq\frac{\delta^{\prime}}{2}f^{\prime\prime}(1).$
Also, the convexity of $f$ implies that $f^{\prime}(x)$ is non-decreasing and $\frac{f(x)}{1-x}$ is non-increasing.
We will prove that for $x<y$, $\{x,y\}\cap(1-\delta,1+\delta)\neq\emptyset$, and $1\not\in[x,y]$,
(A.12) $\exists\alpha_{\delta}>0\quad\textrm{s.t.}\quad\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}.$
When $1-2\delta^{\prime}<x<y<1+2\delta^{\prime}$, (A.12) can be obtained with $\alpha_{\delta_{1}}=\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})^{2}}{4}$ from the following two inequalities
(A.13) $f^{\prime}(y)-f^{\prime}(x)$ $=\int_{x}^{y}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\frac{f^{\prime\prime}(1)}{2}(y-x)$
[…]
$\frac{f(x)}{1-x}-\frac{f(y)}{1-y}$ $=\int_{x}^{y}\frac{-f(\xi)+f^{\prime}(\xi)(\xi-1)}{(1-\xi)^{2}}\mathrm{d}\xi=\int_{x}^{y}\frac{\int_{1}^{\xi}f^{\prime\prime}(z)(z-1)\mathrm{d}z}{(1-\xi)^{2}}\mathrm{d}\xi$
[…]
$=f^{\prime\prime}(1)(y-x)$
$\leq 2(f^{\prime}(y)-f^{\prime}(x)).$
Here in the last inequality, we used (A.13).
When $1<x<1+\delta^{\prime}<1+2\delta^{\prime}<y$, (A.12) can be obtained with $\alpha_{\delta_{2}}=\frac{\delta^{\prime}f^{\prime\prime}(1)}{6}$ from the following two inequalities
(A.15) $f^{\prime}(y)-f^{\prime}(x)$ $=\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1+\delta^{\prime}}^{1+2\delta^{\prime}}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\bigl{(}\frac{1}{x}-\frac{1}{y}\bigr{)},$
$\frac{f(x)}{1-x}-\frac{f(y)}{1-y}$ $=\int_{1}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi-\int_{1}^{x}f^{\prime\prime}(\xi)\frac{x-\xi}{x-1}{\rm d}\xi$
(A.16) $=\frac{y-x}{(y-1)(x-1)}\int_{1}^{x}f^{\prime\prime}(\xi)(\xi-1){\rm d}\xi+\int_{x}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi$
$\leq\frac{1}{x-1}\int_{1}^{x}2f^{\prime\prime}(1)(\xi-1){\rm d}\xi+\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi$
$\leq\delta^{\prime}f^{\prime\prime}(1)+\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi$
$\leq 3\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}.$
Here in the last inequality, we used (A.15).
When $10<x<1-2\delta^{\prime}<1-\delta<y<$ 1, (A.12) can be obtained with $\alpha_{\delta_{3}}=\frac{\alpha}{4}$. In fact, the right hand side of (A.12) can be bounded as
(A.17) $(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}\leq\frac{f(x)}{x},$
where we take $y=1$ in the left hand side to obtain the above inequality. For the left hand side of eq. A.12, using (A.11) leads to
(A.18) $-f^{\prime}(x)$ $=\int_{x}^{1}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1-2\delta^{\prime}}^{1}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\delta^{\prime}f^{\prime\prime}(1)\geq-2f^{\prime}(1-\delta)\geq-2f^{\prime}(y).$
Using Equations A.18, A.17 and A.8 leads to
$\Bigl{(}f^{\prime}(x)-f^{\prime}(y)\Bigr{)}^{2}\geq\frac{f^{\prime}(x)^{2}}{4}\geq\frac{\alpha}{4}\frac{f(x)}{x}\geq\frac{\alpha}{4}(\frac{1}{x}-\frac{1}{y})\bigl{(}\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr{)}.$
Therefore, we can choose the $\delta>0$ defined in eqs. A.10 and A.11. For $x\in(1-\delta,1+\delta)$, $y>0$, $x\neq y$, and $1\not\in[x,y]$ we have (A.12) holds with $\alpha_{\delta}=\min\{\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})}{4},\frac{\delta^{\prime}f^{\prime\prime}(1)}{6},\frac{\alpha}{4}\}$.
For $x\in(1-\delta,1+\delta)$ and $y>0$, the remaining cases are (i) $x=y$ and (ii) $1\in[x,y]$. For (i) (A.12) becomes $0\leq 0$; for (ii) (A.12) follows from 4.2. Hence, (4.4) holds with $\alpha_{\delta}>0$. The proof is complete. ∎
Proof of Lemma 4.3.
We prove that there is a sequence of bounded domains $\{\Omega_{n}\}$ with $\cup_{n=1}^{\infty}\Omega_{n}=\mathbb{R}^{d}$, such that $\int_{\Omega_{n}}\rho\mathrm{d}\theta=\int_{\Omega_{n}}{\rho^{*}}\mathrm{d}\theta$.
Due to the smoothness of $\rho$ and ${\rho^{*}}$, when they are not the same, there exist $\theta_{a}$, $\theta_{b}$, $\delta>0$ and $\epsilon>0$, such that
$\rho(\theta)-{\rho^{*}}(\theta)>\epsilon,\qquad\forall\ \theta\in B_{\delta}(\theta_{a}),$
${\rho^{*}}(\theta)-\rho(\theta)>\epsilon,\qquad\forall\ \theta\in B_{\delta}(\theta_{b}),$
where $B_{r}(\theta)$ is the ball centered at $\theta$ with radius $r$. Let $|B_{r}(\theta)|$ denote the volume of the ball.
We first construct the bounded domain
(A.19) $\Omega_{n}=B_{R_{n}}(0)-\bigl{(}B_{\frac{1}{n}}(\theta_{a})\cup B_{\frac{1}{n}}(\theta_{b})\bigr{)}$
with $n>\frac{1}{\delta}$ and $R_{n}>n$ large enough such that $\int_{B_{R_{n}}^{c}(0)}\rho\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$ and $\int_{B_{R_{n}}^{c}(0)}{\rho^{*}}\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$. Without loss of generality, we assume $\int_{\Omega_{n}}{\rho^{*}}\mathrm{d}\theta>\int_{\Omega_{n}}\rho\mathrm{d}\theta$, then we define
(A.20) $\Omega_{n}(r)=B_{R_{n}}(0)-\bigl{(}B_{r}(\theta_{a})\cup B_{1/n}(\theta_{b})\bigr{)}$
When $r=0$, we have
(A.21) $\int_{{\Omega}_{n}(r)}{\rho^{*}}\mathrm{d}\theta$ $=1-\int_{B_{R_{n}}(0)^{c}}{\rho^{*}}\mathrm{d}\theta-\int_{B_{1/n}(\theta_{b})}{\rho^{*}}\mathrm{d}\theta$
$\leq 1-\int_{B_{1/n}(\theta_{b})}(\rho-\epsilon)\mathrm{d}\theta$
$\leq 1-\int_{B_{1/n}(\theta_{b})}\rho\mathrm{d}\theta-\epsilon|B_{1/n}(0)|$
$<1-\int_{B_{1/n}(\theta_{b})}\rho\mathrm{d}\theta-\int_{B_{R_{n}}(0)^{c}}\rho\mathrm{d}\theta=\int_{{\Omega}_{n}(r)}\rho\mathrm{d}\theta.$
Therefore, by continuity, there exists some $r_{n}\in(0,1/n)$, such that $\int_{{\Omega}_{n}(r_{n})}{\rho^{*}}\mathrm{d}\theta=\int_{{\Omega}_{n}(r_{n})}\rho\mathrm{d}\theta$. Using ${\Omega}_{n}(r_{n})$ to replace $\Omega_{n}$ completes the proof. We can also make $\Omega_{n}$ monotonically increasing by using the union $\cup_{k=1}^{n}\Omega_{k}$ to replace $\Omega_{n}$. ∎
Appendix B Lemmas for Proof of Theorem 5.2
In this section, we provide the proofs of Lemma 5.4 and Lemma 5.5, which are needed in the proof of Theorem 5.2.
Proof of Lemma 5.4.
The convexity of $f$ leads to that
(B.1) $f(y)\geq f(x)+(y-x)f^{\prime}(x),\quad\forall\ y>0,\ x>0.$
Adding $\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}(y-x)$ to both sides gives
$\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}(y-x)\geq(y-x)f^{\prime}(y)-f(y)+f(x)\geq 0,\quad\forall\ y>0,\ x>0.$
Then for $x>0$, we have
(B.2) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(\frac{1}{\mu}-\frac{1}{x})\geq$ $(\frac{1}{\mu}-\frac{1}{x})f^{\prime}(\frac{1}{\mu})-f(\frac{1}{\mu})+f(\frac{1}{x})\geq 0,$
$\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})=$ $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(\frac{1}{\mu}-\frac{1}{x})\mu x(\mu+x)$
$\geq$ $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(\frac{1}{\mu}-\frac{1}{x})\mu^{2}x$
$\geq$ $\bigl{(}(\frac{1}{\mu}-\frac{1}{x})f^{\prime}(\frac{1}{\mu})-f(\frac{1}{\mu})+f(\frac{1}{x})\bigr{)}\mu^{2}x$
(B.3) $=$ $\mu(x-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}x\bigl{(}f(\frac{1}{\mu})-f(\frac{1}{x})\bigr{)}.$
Taking $x={\rho^{*}}/\rho$ in (B.3) leads to
$\Bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})),$
which leads to the desired inequality. This completes the proof. ∎
Proof of Lemma 5.5.
Since $f\in C^{2}$ and $f^{\prime\prime}(1)>0$, there exists $0<\delta_{f}<\frac{1}{4}$, such that for any $x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})$, we have
(B.4) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x).$
For any $x,y$ satisfying $y>x>0$ and $\{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$, we have
(B.5) $f^{\prime}(\frac{1}{x})-f^{\prime}(\frac{1}{y})=\int_{1/y}^{1/x}f^{\prime\prime}(\xi){\rm d}\xi\geq\frac{f^{\prime\prime}(1)}{2}\int_{(\frac{1}{y},\frac{1}{x})\cap(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})}{\rm d}\xi$
$\geq$ $\begin{cases}\frac{f^{\prime\prime}(1)}{2}\Bigl{(}\frac{1}{x}-\frac{1}{y}\Bigr{)}\geq\frac{f^{\prime\prime}(1)}{1+2\delta_{f}}\frac{y-x}{y+x}&1-2\delta_{f}<x<y<1+2\delta_{f}\\ \frac{f^{\prime\prime}(1)}{2}\Bigl{(}\frac{1}{1-2\delta_{f}}-\frac{1}{y}\Bigr{)}\geq\frac{f^{\prime\prime}(1)\delta_{f}}{1-2\delta_{f}}\frac{y-x}{y+x}&0<x\leq 1-2\delta_{f}<1\leq y\leq 1+\delta_{f}\\ \frac{f^{\prime\prime}(1)}{2}\Bigl{(}\frac{1}{x}-\frac{1}{1+2\delta_{f}}\Bigr{)}\geq\frac{f^{\prime\prime}(1)\delta_{f}}{2(1+2\delta_{f})(1+\delta_{f})}\frac{y-x}{y+x}&1\leq x\leq 1+\delta_{f}<1+2\delta_{f}\leq y\end{cases}$
$\geq$ $\alpha_{f}^{\prime}\frac{y-x}{y+x},$
where
$\alpha_{f}^{\prime}=\frac{f^{\prime\prime}(1)\delta_{f}}{2(1+2\delta_{f})(1+\delta_{f})}>0.$
These cases cover all situations with $y>x>0,\ \{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$.
Then we consider (5.9) with $\mu\in[1,1+\delta_{f}]$. When $0<x<\mu$, by using (B.5), we have
(B.6) $\begin{split}\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}\frac{\frac{1}{x}-\frac{1}{\mu}}{\frac{1}{x}+\frac{1}{\mu}}(\mu^{2}-x^{2})=\alpha_{f}^{\prime}(\mu-x)^{2}.\end{split}$
When $\mu<x$, by using (B.5), we have
(B.7) $\begin{split}\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}\frac{\frac{1}{\mu}-\frac{1}{x}}{\frac{1}{x}+\frac{1}{\mu}}(x^{2}-\mu^{2})=\alpha_{f}^{\prime}(\mu-x)^{2}.\end{split}$
If $x=\mu$ then (5.9) becomes $0\geq 0$ and is still true. This completes the proof. ∎